### v1

# Berry-Esséen bound for the Moment Estimation of the fractional Ornstein–Uhlenbeck model under fixed step size discrete observations

Abstract
Let the Ornstein–Uhlenbeck process $\left\{X_{t},\,t\geq 0\right\}$ driven by a fractional Brownian motion $B^{H}$ described by $\mathrm{d}X_{t}=-\theta X_{t}\mathrm{d}t+\mathrm{d}B_{t}^{H},\,X_{0}=0$ with known parameter $H\in(0,\frac{3}{4})$ be observed at discrete time instants $t_{k}=kh,k=1,2,\dots,n$. If $\theta>0$ and if the step size $h>0$ is arbitrarily fixed, we derive Berry-Esséen bound for the ergodic type estimator (or say the moment estimator) $\hat{\theta}_{n}$, i.e., the Kolmogorov distance between the distribution of $\sqrt{n}(\hat{\theta}_{n}-\theta)$ and its limit distribution is bounded by a constant $C_{\theta,H,h}$ times $n^{-\frac{1}{2}}$ and $n^{4H-3}$ when $H\in(0,\,\frac{5}{8}]$ and $H\in(\frac{5}{8},\,\frac{3}{4})$, respectively. This result greatly improve the previous result in literature where $h$ is forced to go zero. Moreover, we extend the Berry-Esséen bound to the Ornstein–Uhlenbeck model driven by a lot of Gaussian noises such as the sub-bifractional Brownian motion and others. A few ideas of the present paper come from Haress and Hu (2021), Sottinen and Viitasaari (2018), and Chen and Zhou (2021).
keywords:
Fractional Brownian motion; Fourth moment theorem; Berry-Esséen bound; Fractional Ornstein–Uhlenbeck; Sub-bifractional Brownian motion; Fractional Gaussian process; Kolmogorov distance; Malliavin calculus.
{amscode}
62M09,60G22,60H10,60H30
1 Introduction and main results
The fractional Ornstein-Uhlenbeck processes $\left\{X_{t}:t\geq 0\right\}$ is known as the solution of the Langevin equation
$\mathrm{d}X_{t}=-\theta X_{t}\mathrm{d}t+\sigma\mathrm{d}B_{t}^{H},\quad t\in[0,T],\quad$ (1)
where $\sigma>0,\,\theta>0$ are the unknown parameter and $B_{t}^{H}$ is the fractional Brownian motion (fBm) with Hurst index $H\in(0,1)$. The problem of estimation of part or all parameters $(\theta,\sigma,H)$ has been intensively studied in the last decades, see for example [9, 8] and the references therein. In the literature, it is often assume that the process $\left\{X_{t}:t\geq 0\right\}$ is observed continuously or discretely with a step size $h_{n}$ which is forced to go zero as $n\to\infty$, see [11] for example.
Recently, the assumption of the step size $h_{n}\to 0$ as $n\to\infty$ is finally removed in [14]. In detail, denote $\{X_{jh}:j=1,\cdots,n\}$ the discrete-time observations of the processes $\{X_{t}:t\geq 0\}$, sampled at equidistant time points $t_{j}=jh$ with fixed step size $h$ and $n$ is the sample size. They propose an ergodic type statistical estimator $(\hat{\theta}_{n},\,\hat{H}_{n},\hat{\sigma}_{n})$ for all the parameter $(\theta,H,\sigma)$ and show the strong consistence and the central limit theorem.
Let us recall the moment estimator for unknown parameter $\theta$ under discrete observations $\{X_{jh}:j=1,\cdots,n\}$:
$\hat{\theta}_{n}=\left(\frac{1}{{H}\Gamma(2H)n}\sum_{j=1}^{n}X_{jh}^{2}\right)^{-\frac{1}{2H}}.$ (2)
In the present paper, we aim to show, under the framework of [14], the rate of convergence of the estimator $\hat{\theta}_{n}$ in the Kolmogorov distance, which is called the Berry-Esséen type upper bound in literature [6]. We point out that in [15], the rate of convergence for the estimator is obtained in the $p$-Wasserstein distance.
For simplicity, we assume that $X_{0}=0$ and that $H,\,\sigma$ are known and $\sigma=1$ in (1) from now on. Other initial value of $X_{0}$ and other parameter value of $\sigma$ can be treated exactly in the same way.
We emphasize again that all the previous result concerning the Berry-Esséen type upper bound for the parameters estimate problem of the fractional Ornstein-Uhlenbeck process is under the assumption of continuous observations or discrete observations with the step size $h_{n}\to 0$ as $n\to\infty$, see [24, 6, 11] for example. The first contribution of the present paper is to derived the upper bound of the Kolmogorov distance between $\sqrt{n}(\hat{\theta}_{n}-\theta)$ and its limit distribution. We state it as follows.
Theorem 1.1.
Assume that $H\in(0,\frac{3}{4})$ and the fractional Ornstein-Uhlenbeck process $\{X_{t}:t\geq 0\}$ is defined as in (1). If the process is observed at discrete time instants $t_{k}=kh,\,k=1,2,\dots,n$ and the estimator $\hat{\theta}_{n}$ is given by (2), then there exists a positive constant $C_{\theta,H,h}$ independent of $n$ such that when $n$ is sufficiently large,
$\begin{split}d_{Kol}\left(\sqrt{n}(\hat{\theta}_{n}-\theta),\mathcal{N}\right)\leq C_{\theta,H,h}\times\left\{\begin{array}[]{ll}\frac{1}{\sqrt{n}},&\quad\text{if }H\in(0,\,\frac{5}{8}],\\ \frac{1}{n^{3-4H}},&\quad\text{if }H\in(\frac{5}{8},\,\frac{3}{4}),\end{array}\right.\end{split}$ (3)
where the normal random variable $\mathcal{N}\sim N(0,\sigma_{1}^{2})$ with $\sigma_{1}^{2}=\frac{\theta^{2}\sigma_{B}^{2}}{4H^{2}a^{2}}$ and $a=H\Gamma(2H)\theta^{-2H}$, $\sigma_{B}^{2}=2\sum_{j=-\infty}^{+\infty}\rho_{0}^{2}(jh)<+\infty$.
Remark 1.
When $H=\frac{3}{4}$, the scaling before $(\hat{\theta}_{n}-\theta)$ is $\sqrt{\frac{n}{\log n}}$ and the corresponding upper bound is $\frac{1}{\log n}$. However, if $H>\frac{3}{4}$, the central limit theorem about the convergence of parameter will be no longer satisfy. We refer the reader to Hu et al. [16], Haress and Hu [14] and Chen et al. [6] for details.
The assumption of the step size $h_{n}\to 0$ as $n\to\infty$ in previous literature [11] is due to their method by which the result of the discrete observation is transitioned from that of the continuous observation. The idea of [14] is to deal with the double Wiener chaos random variable $W_{n}$ (see below) concerning the discrete observation directly. Our proof follows this idea.
The second aim of the present paper is to extend Theorem 1.1 to the Ornstein-Uhlenbeck models driven by some well-known Gaussian noise such as the sub-fractional Brownian motion, the bi-fractional Brownian motion, and the sub-bifractional Brownian motion. Now, let us recall these fractional Gaussian processes firstly.
Example 1.2.
The sub-fractional Brownian motion $\{S^{H}(t),t\geq 0\}$ with parameter $H\in(0,1)$ has the covariance function
$R(t,s)=s^{2H}+t^{2H}-\frac{1}{2}\left((s+t)^{2H}+|t-s|^{2H}\right).$
Example 1.3.
The bi-fractional Brownian motion $\{B^{H^{\prime},K}(t),t\geq 0\}$ with parameters $H^{\prime}\in(0,1),K\in(0,2)$ and $H:=H^{\prime}K\in(0,1)$ has the covariance function
$R(t,s)=\frac{1}{2^{K}}\left((s^{2H^{\prime}}+t^{2H^{\prime}})^{K}-|t-s|^{2H^{\prime}K}\right).$
Example 1.4.
The covariance function of the generalized sub-fractional Brownian motion (also known as the sub-bifractional Brownian motion), $S^{H^{\prime},K}(t)$, with parameters $H^{\prime}\in(0,1)$ and $K\in(0,2)$, such that $H:=H^{\prime}K\in(0,1)$, is given by:
$R(s,\,t)=(s^{2H^{\prime}}+t^{2H^{\prime}})^{K}-\frac{1}{2}\left[(t+s)^{2H^{\prime}K}+\left|t-s\right|^{2H^{\prime}K}\right].$
When $K=1$, it degenerates to the sub-fractional Brownian motion $S^{H}(t)$. Some properties of the process for $K\in(0,1)$ and $K\in(1,2)$ have been studied in [12, 23].
Example 1.5.
The generalized fractional Brownian motion is an extension of both fBm and sub-fractional Brownian motion. Its covariance function is given by:
$R(s,\,t)=\frac{(a+b)^{2}}{2(a^{2}+b^{2})}(s^{2H}+t^{2H})-\frac{ab}{a^{2}+b^{2}}(s+t)^{2H}-\frac{1}{2}\left|t-s\right|^{2H},$
where $H\in(0,1)$ and $(a,b)\neq(0,0)$ (see [26]).
The strategy we will use is not to derive the upper Berry-Esséen bound for the Ornstein-Uhlenbeck models driven by the above four types of Gaussian noises one by one. We will use a more general condition in terms of the covariance functions of the Gaussian noise cited from [9, 8] and show that the desired upper Berry-Esséen bound holds for all the Ornstein-Uhlenbeck models driven by the type of general Gaussian noise.
Let us rewrite the Ornstein-Uhlenbeck model as $(Z_{t})_{t\in[0,T]}$, which is the solution of the Langevin equation
$\mathrm{d}Z_{t}=-\theta Z_{t}\mathrm{d}t+\mathrm{d}G_{t},\quad t\in[0,T],\quad Z_{0}=0$ (4)
where the driving Gaussian noise $G_{t}$ satisfies the following hypothesis.
HYPOTHESIS 1.6.
For $H\in(0,1)$ and $H\neq\frac{1}{2}$, the covariance function $R(s,t)=\mathbb{E}[G_{t}G_{s}]$ of the centered Gaussian process $(G_{t})_{t\in[0,T]}$ with $G_{0}=0$ satisfies the following three hypotheses:
  • ($H_{1}$)
    For any fixed $s\in[0,T]$, $R(s,t)$ is an absolutely function with respect to $t$ on interval $[0,T]$.
  • ($H_{2}$)
    For any fixed $t\in[0,T]$, the difference
    $\begin{split}\frac{\partial R(s,t)}{\partial t}-\frac{\partial R^{B}(s,t)}{\partial t}\end{split}$ (5)
    is an absolutely continuous function with respect to $s\in[0,T]$, where $R^{B}(s,t)$ is the covariance function of fBm $(B^{H}_{t})_{t\in[0,T]}$.
  • ($H_{3}$)
    There exists a positive constant $C$ independent of $T$ such that
    $\begin{split}\bigg{|}\frac{\partial}{\partial s}\bigg{(}\frac{\partial R(s,t)}{\partial t}-\frac{\partial R^{B}(s,t)}{\partial t}\bigg{)}\bigg{|}\leq C(ts)^{H-1},\end{split}$ (6)
    holds.
It is clear that the four Gaussian noises from Example 1.2 to Example 1.5 satisfy Hypothesis 1.6, see [9, 8]. Now we give our second contribution of the present paper.
Theorem 1.7.
Assume that the Ornstein-Uhlenbeck model $\{Z_{t}:t\geq 0\}$ is defined as in (4) and that the process is observed at discrete time instants $t_{k}=kh,\,k=1,2,\dots,n$ and the estimator $\hat{\theta}_{n}$ is given by
$\hat{\theta}_{n}=\left(\frac{1}{{H}\Gamma(2H)n}\sum_{j=1}^{n}Z_{jh}^{2}\right)^{-\frac{1}{2H}}.$ (7)
If the driving noise satisfies Hypothesis 1.6 with $H\in(0,\frac{1}{2})$, then there exists a positive constant $C_{\theta,H,h}$ independent of $n$ such that when $n$ is sufficiently large,
$d_{Kol}\left(\sqrt{n}(\hat{\theta}_{n}-\theta),\mathcal{N}\right)\leq C_{\theta,H,h}\times\frac{1}{\sqrt{n}}$ (8)
where $\mathcal{N}$ is the same as in Theorem 1.1.
The inequality (6) is good enough in some applications, see [9, 8]. However, in some situations, a more steep inequality (9) is needed, see [5]. We write it as a new Hypothesis.
HYPOTHESIS 1.8.
For $H\in(0,1)$ and $H\neq\frac{1}{2}$, the covariance function $R(s,t)=\mathbb{E}[G_{t}G_{s}]$ of the centered Gaussian process $(G_{t})_{t\in[0,T]}$ with $G_{0}=0$ satisfies the above ($H_{1}$), ($H_{2}$) and the following:
  • ($H_{3}^{\prime}$)
    There exists a positive constants $C_{1},\,C_{2}$ which depend only on $H^{\prime},\,K$ such that the inequality
    $\left|\frac{\partial}{\partial s}\left(\frac{\partial R(s,t)}{\partial t}-\frac{\partial R^{B}(s,t)}{\partial t}\right)\right|\leq C_{1}(t+s)^{2H-2}+C_{2}(s^{2H^{\prime}}+t^{2H^{\prime}})^{K-2}(st)^{2H^{\prime}-1}$ (9)
    holds, where $H^{\prime}\in(\frac{1}{2},1),\,K\in(0,2)$ and $H:=H^{\prime}K\in(0,1)$.
Clearly, both Example 1.2 and Example 1.5 satisfy Hypothesis 1.8. An additional requirement $H^{\prime}\in(\frac{1}{2},1)$ for both Example 1.4 and Example 1.3 makes Hypothesis 1.8 hold.
Theorem 1.9.
Assume that both the Ornstein-Uhlenbeck model $\{Z_{t}:t\geq 0\}$ and the estimator $\hat{\theta}_{n}$ are given as in Theorem 1.7. If the driving noise satisfies Hypothesis 1.8 with $H\in(\frac{1}{2},\frac{3}{4})$, then there exists a positive constant $C_{\theta,H,h}$ independent of $n$ such that when $n$ is sufficiently large,
$d_{Kol}\left(\sqrt{n}(\hat{\theta}_{n}-\theta),\mathcal{N}\right)\leq C_{\theta,H,h}\times\left\{\begin{array}[]{ll}\frac{1}{\sqrt{n}},&\quad\text{if }H\in(\frac{1}{2},\,\frac{5}{8}],\\ \frac{1}{n^{3-4H}},&\quad\text{if }H\in(\frac{5}{8},\,\frac{3}{4}),\end{array}\right.$ (10)
where $\mathcal{N}$ is the same as in Theorem 1.1.
Remark 2.
The assumption of $H^{\prime}\in(\frac{1}{2},1)$ in Hypothesis ($H_{3}^{\prime}$) rules out $H^{\prime}\in(0,\frac{1}{2}]$, which is not an essential but only a technical requirement.
Based on Theorem 1.7 and Theorem 1.9, we can finish the second aim of the present paper as follows.
Corollary 1.10.
Assume that the Ornstein-Uhlenbeck model $\{Z_{t}:t\geq 0\}$ is driven by the sub-fractional Brownian motion, the bi-fractional Brownian motion, the sub-bifractional Brownian motion or the generalized fractional Brownian motion, and that the estimator $\hat{\theta}_{n}$ is given as in Theorem 1.7. If an additional requirement $H^{\prime}\in(\frac{1}{2},1)$ holds for both the bi-fractional Brownian motion and the sub-bifractional Brownian motion, then there exists a positive constant $C_{\theta,H,h}$ independent of $n$ such that when $n$ is sufficiently large,
$d_{Kol}\left(\sqrt{n}(\hat{\theta}_{n}-\theta),\mathcal{N}\right)\leq C_{\theta,H,h}\times\left\{\begin{array}[]{ll}\frac{1}{\sqrt{n}},&\quad\text{if }H\in(0,\,\frac{5}{8}],\\ \frac{1}{n^{3-4H}},&\quad\text{if }H\in(\frac{5}{8},\,\frac{3}{4}),\end{array}\right.$ (11)
where $\mathcal{N}$ is the same as in Theorem 1.1.
The paper is organized as follows. In Section 2, we recall some known results of stochastic analysis. Proof of Theorem 1.1 is given in Section 3. Proof of Theorem 1.7 and Theorem 1.9 are given in Section 4. To make the paper more readable, we delay some technical calculations in Appendix.
2 Preliminaries
This section provides a concise overview of foundational elements about Gaussian stochastic analysis and the Berry-Esséen type upper bound quantifying the distance of two normal random variables. Given a complete probability space $(\Omega,\mathcal{F},P)$, we denote by $\{G_{t}:t\in[0,T]\}$ a continuous centered Gaussian process on this space with covariance function
$\mathbb{E}(G_{t}G_{s})=R(t,s),\quad s,t\in[0,T].$
Let $\mathfrak{H}$ be the associated reproducing kernel Hilbert space of the Gaussian process $G$, which is defined as the closure of the space of all real-valued step functions on $[0,T]$, equipped with the inner product
[…]
for any $0\leq a<b\leq T$ and $0\leq c<d\leq T$. Denote $\{G(h):h\in\mathfrak{H}\}$ by the isonormal Gaussian process on the above probability space $(\Omega,\mathcal{F},P)$ with following representation
$G(h)=\int_{[0,T]}h(t)\mathrm{d}G_{t},\ \forall\ h\in\mathfrak{H},$ (1)
which is indexed by the elements in the Hilbert space $\mathfrak{H}$ and satisfies Itô’s isometry:
$\mathbb{E}\left[G(g)G(h)\right]=\langle g,h\rangle_{\mathfrak{H}},\quad\forall\ g,h\in\mathfrak{H}.$ (2)
The key point lies in establishing the explicit formulas for the inner product in the Hilbert space $\mathfrak{H}$, which follow the idea of [9, 7]. To elaborate it, we first define the covariance function of fBm $B^{H}$ by $R^{B}(s,t)=\mathbb{E}[B^{H}_{s}B^{H}_{t}]$, and subsequently denote the associated canonical Hilbert space by $\mathfrak{H}_{1}$ throughout the paper. When $H\in(\frac{1}{2},1)$ or the Lebesgue measure of intersection of the supports about two function $f,g\in\mathfrak{H}$ is zero, Mishura [19] provides that
[…]
Suppose that $\mathcal{V}_{[0,T]}$ is the set of functions of bounded variation in $[0,T]$, and by $\mathcal{B}([0,T])$ the Borel $\sigma$-algebra on $[0,T]$. When $H\in(0,\frac{1}{2})$, for any two functions in the set $\mathcal{V}_{[0,T]}$, Chen et al. [8, 1] propose a new inner product in the Hilbert space $\mathfrak{H}_{1}$ as following,
$\langle f,\,g\rangle_{\mathfrak{H}_{1}}$ $=H\int_{[0,T]^{2}}f(t)\left|t-s\right|^{2H-1}\mathrm{sgn}(t-s)\mathrm{d}t\nu_{g}(\mathrm{d}s),\quad\forall f,\,g\in\mathcal{V}_{[0,T]},$ (3)
where $\nu_{g}(\mathrm{d}s):=\mathrm{d}\nu_{g}(s)$, and $\nu_{g}$ is the restriction on $\left({[0,T]},\mathcal{B}({[0,T]})\right)$ of the signed Lebesgue-Stieljes measure $\mu_{g^{0}}$ on $\left(\mathbb{R},\mathcal{B}(\mathbb{R})\right)$, where $g^{0}(x)$ is defined by
$g^{0}(x)=\left\{\begin{array}[]{ll}g(x),&\quad\text{if}~{}x\in[0,T],\\ 0,&\quad\text{otherwise}.\end{array}\right.$
Furthermore, if $g^{\prime}(\cdot)$ is interpreted as the distributional derivative of $g(\cdot)$, the formula (3) admits the following representation:
$\langle f,\,g\rangle_{\mathfrak{H}_{1}}$ $=H\int_{[0,T]^{2}}f(t)g^{\prime}(s)\left|t-s\right|^{2H-1}\mathrm{sgn}(t-s)\mathrm{d}t\mathrm{d}s,\quad\forall f,\,g\in\mathcal{V}_{[0,T]}.$ (4)
Next, for the general Gaussian process $G$ and the associated reproducing kernel Hilbert space $\mathfrak{H}$, if any two functions $f,\,g\in\mathcal{V}_{[0,T]}$, Jolis [17] gives a inner product formula in Theorem 2.3 as following,
$\langle f,\,g\rangle_{\mathfrak{H}}=\int_{[0,T]^{2}}R(s,t)\mathrm{d}\big{(}\nu_{f}\times\nu_{g}\big{)}(s,t),$ (5)
where $\nu_{g}$ is same as in equation (3). Then, under Hypotheses $(H_{1})$-$(H_{2})$, the relationship between the inner products of two functions in the Hilbert spaces $\mathfrak{H}$ and $\mathfrak{H}_{1}$ satisfies that
$\langle f,\,g\rangle_{\mathfrak{H}}-\langle f,\,g\rangle_{\mathfrak{H}_{1}}=\int_{0}^{T}f(t)\mathrm{d}t\int_{0}^{T}g(s)\frac{\partial}{\partial s}\left(\frac{\partial R(s,t)}{\partial t}-\frac{\partial R^{B}(s,t)}{\partial t}\right)\mathrm{d}s.$ (6)
Moreover, when the intersection of these two functions’ supports is of Lebesgue measure zero, we have
$\langle f,\,g\rangle_{\mathfrak{H}}=\int_{[0,T]^{2}}f(t)g(s)\frac{\partial^{2}R(t,s)}{\partial t\partial s}\mathrm{d}t\mathrm{d}s.$ (7)
Finally, we introduce the Berry–Esséen bounds (so-called “Stein’s method”) estimating the distance between two probability distributions. Recall that the Kolmogorov distance between two random variables $\xi,\eta$ as
$d_{Kol}(\xi,\eta):=\sup_{z\in\mathbb{R}}\left|P\left(\xi\leq z\right)-P\left(\eta\leq z\right)\right|.$
Let the function
$y=f(x)=\left(\frac{1}{{H}\Gamma(2H)}x\right)^{-\frac{1}{2H}}.$ (8)
Its inversion function is
$x:=g(y)=f^{-1}(y)=H\Gamma(2H)y^{-2H}.$ (9)
If $X\geq 0$ almost surely, the following lemma provides an estimate of the Kolmogorov distance between the random variable $f(X)$ and one normal random variable by means of that between the random variable $X$ and another normal random variable (see [9, 8, 24].)
Lemma 2.1.
Let $T$ be any positive real number and $\xi\sim N(0,\sigma^{2}_{1})$ and $\eta\sim N(0,\sigma^{2}_{2})$ and the two functions $f$ and $g$ given by (8) and (9), respectively. If a random variable $X\geq 0$ almost surely, then there exists a positive constant $C$ independent of $T$ such that
$d_{Kol}(\sqrt{T}(f(X)-\theta),\,\xi)\leq C\times\left(d_{Kol}(\sqrt{T}(X-\mathbb{E}[X]),\eta)+\sqrt{T}\left|\mathbb{E}[X]-g(\theta)\right|+\frac{1}{\sqrt{T}}\right),$
where the two variance $\sigma^{2}_{2},\,\sigma^{2}_{1}$ satisfy the following relation:
$\sigma^{2}_{2}=\big{(}g^{\prime}(\theta)\big{)}^{2}\times{\sigma^{2}_{1}}.$ (10)
The relation (10) comes from the delta method, please refer to chapter 3 of [25]. We point that in the previous literature [9, 8], the random variable $X$ is taken as
$\frac{1}{T}\int_{0}^{T}Z_{t}^{2}\mathrm{d}t,$
however, in the present paper, we take $T=n$ and take the random variable $X$ as
$\frac{1}{n}\sum_{j=1}^{n}X_{jh}^{2};\text{ and }\,\,\frac{1}{n}\sum_{j=1}^{n}Z_{jh}^{2},$ (11)
where $\{X_{t}:t\geq 0\},\{Z_{t}:t\geq 0\}$ are the Ornstein-Uhlenbeck model defined as in (1) and (4), respectively.
3 Proof of Theorem 1.1
In this section, the Ornstein-Uhlenbeck model is defined as in (1). By Lemma 2.1, we need to study the property of the second moment of sample path for the fractional Ornstein-Uhlenbeck process defined as in (11). It is convenient to introduce a new notation and rewrite it as follows:
$B_{n}:=\frac{1}{n}\sum_{j=1}^{n}X_{jh}^{2}.$ (12)
Next, we will elaborate the limit of $\mathbb{E}(B_{n})$ as $n$ large enough and its convergence rate.
Proposition 3.1.
Let $H\in(0,1)$ and $B_{n}$ be defined as in (12). When $n$ large enough, the exist a constant $C$ independent of $n$ such that
$\left|\mathbb{E}(B_{n})-a\right|\leq C\times\frac{1}{n},$ (13)
where the constant $a=g(\theta)=H\Gamma(2H)\theta^{-2H}$.
Proof.
Through standard computations, the fractional Ornstein-Uhlenbeck processes $X_{t}$, known as the solution of (1), admits the explicit representation:
$X_{t}=\int_{0}^{t}e^{-\theta(t-s)}\mathrm{d}B_{s}^{H},\quad t\geq 0.$ (14)
Furthermore, Let $\{Y_{t},t\in\mathbb{R}\}$ represent the stationary solution of fractional Ornstein-Uhlenbeck processes, expressed as
$Y_{t}=\int_{-\infty}^{t}e^{-\theta(t-s)}\mathrm{d}B_{s}^{H},\quad t\in\mathbb{R}.$ (15)
The stationary property of $Y_{t}$ ensures that
$\mathbb{E}\left(Y_{t}^{2}\right)=\mathbb{E}\left(Y_{0}^{2}\right)=a,$ (16)
where the last equality is from Lemma 19 in Hu et al. [16]. Consequently, by the definition of $B_{n}$ we have
$\left|\mathbb{E}(B_{n})-a\right|=\left|\frac{1}{n}\sum_{j=1}^{n}\left(\mathbb{E}\left(X_{jh}^{2}\right)-\mathbb{E}\left(Y_{jh}^{2}\right)\right)\right|\leq\frac{1}{n}\sum_{j=1}^{n}\left(\mathbb{E}\left|X_{jh}^{2}-Y_{jh}^{2}\right|\right).$ (17)
Crucially, $X_{t}$ and $Y_{t}$ satisfy the relationship
$X_{t}=Y_{t}-e^{-\theta t}Y_{0},\quad\forall t\geq 0.$ (18)
Then the Cauchy-Schwarz inequality and triangle inequality yield
$\left|\mathbb{E}\left(X_{t}^{2}-Y_{t}^{2}\right)\right|=e^{-\theta t}\left|\mathbb{E}\left(Y_{0}\left(e^{-\theta t}Y_{0}-2Y_{t}\right)\right)\right|\leq 3ae^{-\theta t}.$ (19)
Substituting this result into (17), we obtain the desired result. ∎
3.1 The second moment and cumulants of the random variable $W_{n}$
To establish Theorem 1.1, it is necessary to derive the Berry-Esséen bound for the random variable $W_{n}$ based on the idea of [13, 9, 8, 24], which is a second Wiener chaos with respect to the fBm $B_{t}^{H}$ with the form
$W_{n}:=\sqrt{n}\left(B_{n}-\mathbb{E}(B_{n})\right)=\frac{1}{\sqrt{n}}\sum_{j=1}^{n}\left(X_{jh}^{2}-\mathbb{E}\left(X_{jh}^{2}\right)\right).$ (20)
Guided by the optimal fourth moment theorem, our analysis focuses on: 1. Estimating the limit and convergence rate of the second moment of $W_{n}$; 2. Establishing upper bounds for its third and fourth cumulants. These objectives are expounded in the following two propositions.
Proposition 3.2.
Let $H\in(0,\frac{3}{4})$ and $W_{n}$ be defined as in (20). When $n$ is large enough, the exist a constant $C$ independent of $n$ such that
$\begin{split}\left|\mathbb{E}(W_{n}^{2})-\sigma_{B}^{2}\right|\leq C\times\left\{\begin{array}[]{ll}\frac{1}{n},&\quad\text{if }H\in(0,\,\frac{1}{2}],\\ \frac{1}{n^{3-4H}},&\quad\text{if }H\in(\frac{1}{2},\,\frac{3}{4}),\end{array}\right.\end{split}$ (21)
where $\sigma_{B}^{2}$ is a series given by
$\begin{split}\sigma_{B}^{2}=2\sum_{j=-\infty}^{+\infty}\rho_{0}^{2}(jh)<+\infty.\end{split}$ (22)
Proof.
Firstly, we derive the convergence of above series $\sigma_{B}^{2}$ and denote
$\begin{split}\rho(t,s)=\mathbb{E}(X_{t}X_{s}),\quad\rho_{0}(t,s)=\mathbb{E}(Y_{t}Y_{s})\end{split}$ (23)
by the covariance function of fractional Ornstein-Uhlenbeck processes $X_{t}$ and that of stationary process $Y_{t}$. Moreover, due to the stationary property of $Y_{t}$, we can write it’s covariance function as
$\begin{split}\rho_{0}(|t-s|)=\rho_{0}(t,s)=\mathbb{E}(Y_{t}Y_{s}),\quad\forall s,t\in\mathbb{R}.\end{split}$ (24)
Specially, $\rho_{0}(t)=\mathbb{E}(Y_{t}Y_{0})$. From Theorem 2.3 of Cheridito et al. [10], we know that when $t$ is large enough,
$\begin{split}\left|\rho_{0}(t)\right|=O(|t|^{2H-2}),\end{split}$ (25)
which implies that the series $\sigma_{B}^{2}$ converges, i.e.,
$\begin{split}\sigma_{B}^{2}=2\sum_{j=-\infty}^{+\infty}\rho_{0}^{2}(jh)<+\infty\end{split}$ (26)
if and only if $0<H<\frac{3}{4}$, please also refer to Lemma 6.3 of Nourdin [20].
Secondly, according to the product formula of Wiener-Itô multiple integrals, the second moment of second Wiener chaos $W_{n}$ can be rewritten as
$\begin{split}\mathbb{E}(W_{n}^{2})=\frac{2}{n}\sum_{j,l=1}^{n}\rho^{2}(jh,lh).\end{split}$ (27)
Then, the triangle inequlity implies that
$\begin{split}\left|\mathbb{E}(W_{n}^{2})-\sigma_{B}^{2}\right|&\leq\frac{2}{n}\sum_{j,l=1}^{n}\left|\rho^{2}(jh,lh)-\rho_{0}^{2}(jh,lh)\right|+\left|\frac{2}{n}\sum_{j,l=1}^{n}\rho_{0}^{2}(\left|j-l\right|h)-\sigma_{B}^{2}\right|\\ &:=S_{1}+S_{2}.\end{split}$ (28)
For the term $S_{1}$, the fact $\mathbb{E}\left(Y_{t}^{2}\right)=\mathbb{E}\left(Y_{0}^{2}\right)=a$ and $\sup_{t\geq 0}\mathbb{E}\left(X_{t}^{2}\right)<\infty$ (see Theorem 3.1 of Balde et al. [2] ) and Cauchy-Schwarz inequality imply that
$\begin{split}\left|\rho^{2}(jh,lh)-\rho_{0}^{2}(jh,lh)\right|&=\left|\left(\rho(jh,lh)+\rho_{0}(jh,lh)\right)\cdot\left(\rho(jh,lh)-\rho_{0}(jh,lh)\right)\right|\\ &\leq C\left|\rho(jh,lh)-\rho_{0}(jh,lh)\right|\end{split}$ (29)
Combining the relationship (18) and a well-known fact (see Theorem 2.3 of Cheridito et al. [10]) as following
$\begin{split}\left|\rho_{0}(t-s)\right|\leq C(1+|t-s|)^{2H-2},\end{split}$ (30)
we have
$\begin{split}\left|\rho(t,s)-\rho_{0}(t,s)\right|&=\left|\mathbb{E}\left[\left(Y_{t}-e^{-\theta t}Y_{0}\right)\left(Y_{s}-e^{-\theta s}Y_{0}\right)\right]-\mathbb{E}\left(Y_{t}Y_{s}\right)\right|\\ &=\left|e^{-\theta(t+s)}\mathbb{E}\left(Y_{0}^{2}\right)-e^{-\theta t}\mathbb{E}\left(Y_{s}Y_{0}\right)-e^{-\theta s}\mathbb{E}\left(Y_{t}Y_{0}\right)\right|\\ &\leq C\left[e^{-\theta(t+s)}+e^{-\theta t}\left(1+\left|s\right|\right)^{2H-2}+e^{-\theta s}\left(1+\left|t\right|\right)^{2H-2}\right].\end{split}$ (31)
Substituting this estimation into (29) yields
$\begin{split}\left|\rho^{2}(jh,lh)-\rho_{0}^{2}(jh,lh)\right|\leq C\left[e^{-\theta h(l+j)}+e^{-\theta jh}\left(1+l\right)^{2H-2}+e^{-\theta lh}\left(1+j\right)^{2H-2}\right].\end{split}$ (32)
Consequently, we obtain
$\begin{split}S_{1}&\leq\frac{C}{n}\left[\sum_{j,l=1}^{n}e^{-\theta h(l+j)}+\sum_{j,l=1}^{n}e^{-\theta jh}\left(1+l\right)^{2H-2}\right]\\ &\leq\frac{C}{n}\left[\int_{1}^{\infty}\int_{1}^{\infty}e^{-\theta h(x+y)}\mathrm{d}x\mathrm{d}y+\int_{1}^{\infty}e^{-\theta hx}\mathrm{d}x\int_{1}^{n}y^{2H-2}\mathrm{d}y\right]\\ &\leq\frac{C}{n}\left[1+n^{(2H-1)\vee 0}\right]\leq C\times\left\{\begin{array}[]{ll}\frac{1}{n},&\quad\text{if }H\in(0,\,\frac{1}{2}],\\ \frac{1}{n^{2(1-H)}},&\quad\text{if }H\in(\frac{1}{2},\,1).\end{array}\right.\end{split}$ (33)
For the term $S_{2}$, we firstly have known that if and only if $0<H<\frac{3}{4}$,
$\begin{split}\sigma_{B}^{2}=2\sum_{j=-\infty}^{+\infty}\rho_{0}^{2}(jh)<\infty.\end{split}$ (34)
And then, making the change of variable $k=j-l$ yields
$\begin{split}\frac{1}{n}\sum_{j,l=1}^{n}\rho_{0}^{2}(\left|j-l\right|h)=\sum_{k=1-n}^{n-1}\rho_{0}^{2}(\left|k\right|h)(1-\frac{|k|}{n})=\sum_{k=1-n}^{n-1}\rho_{0}^{2}(\left|k\right|h)-\frac{1}{n}\sum_{k=1-n}^{n-1}\rho_{0}^{2}(\left|k\right|h)|k|.\end{split}$ (35)
Therefore, we can scale $S_{2}$ as following
$\begin{split}S_{2}=\left|\frac{2}{n}\sum_{j,l=1}^{n}\rho_{0}^{2}(\left|j-l\right|h)-\sigma_{B}^{2}\right|\leq 2\sum_{|k|\geq n}^{\infty}\rho_{0}^{2}(\left|k\right|h)+\frac{4}{n}\sum_{k=1}^{n}k\rho_{0}^{2}(kh).\end{split}$ (36)
Since the inequality (30) implies $\left|\rho_{0}(kh)\right|\leq C(1+k)^{2H-2}$, then for $0<H<\frac{3}{4}$ we have
$\begin{split}\sum_{|k|\geq n}^{\infty}\rho_{0}^{2}(\left|k\right|h)\leq C\sum_{k=n+1}^{\infty}(1+|k|)^{2(2H-2)}\leq C\int_{n}^{\infty}x^{2(2H-2)}\mathrm{d}x\leq Cn^{4H-3},\end{split}$ (37)
$\begin{split}\frac{4}{n}\sum_{k=1}^{n}k\rho_{0}^{2}(kh)&\leq\frac{C}{n}\sum_{k=1}^{n}(1+k)^{2(2H-2)+1}\leq\frac{C}{n}\int_{n}^{\infty}x^{2(2H-2)+1}\mathrm{d}x\\ &\leq Cn^{(4H-2)\vee 0-1}=C\times\left\{\begin{array}[]{ll}\frac{1}{n},&\quad\text{if }H\in(0,\,\frac{1}{2}],\\ \frac{1}{n^{1-2H}},&\quad\text{if }H\in(\frac{1}{2},\,\frac{3}{4}).\end{array}\right.\end{split}$ (38)
As a result, we get the estimation of $S_{2}$ as
$\begin{split}S_{2}\leq C\times\left\{\begin{array}[]{ll}\frac{1}{n},&\quad\text{if }H\in(0,\,\frac{1}{2}],\\ \frac{1}{n^{1-2H}},&\quad\text{if }H\in(\frac{1}{2},\,\frac{3}{4}).\end{array}\right.\end{split}$ (39)
Substituting the estimations (33) and (39) into (28) with the fact that $32(H-1)<4H-$ 3 if $H\in(\frac{1}{2},\,\frac{3}{4})$, we obtain the desired result. ∎
Next, we derive the the upper bounds of the third and fourth cumulants of $W_{n}$.
Proposition 3.3.
Let $H\in(0,\frac{3}{4})$ and $W_{n}$ be defined as in (20). Then for large enough $n$, we have
$\begin{split}\max\left\{\left|k_{3}(W_{n})\right|,k_{4}(W_{n})\right\}\leq C\times\left\{\begin{array}[]{ll}\frac{1}{\sqrt{n}},&\quad\text{if }H\in(0,\,\frac{2}{3}],\\ n^{\frac{3}{2}(4H-3)},&\quad\text{if }H\in(\frac{2}{3},\,\frac{3}{4}).\end{array}\right.\end{split}$ (40)
Proof.
The core approach involves comparing the third and fourth cumulants of $W_{n}$ with those of $\overline{W}_{n}$, which denote by a random variable
$\overline{W}_{n}=\frac{1}{\sqrt{n}}\sum_{j=1}^{n}\left(Y_{jh}^{2}-\mathbb{E}\left(Y_{jh}^{2}\right)\right),$ (41)
It is clear that $\overline{W}_{n}$ also belongs to the second Wiener chaos with respect to fBm. Then, we apply the product formula of Wiener-Itô multiple integrals to compute the cumulants of second Wiener chaos $W_{n}$ as following
$\begin{split}k_{3}(W_{n}):=\mathbb{E}(W_{n}^{3})=\frac{8}{n^{3/2}}\sum_{j,k,l=1}^{n}\rho(jh,kh)\rho(kh,lh)\rho(lh,jh),\end{split}$ (42)
$\begin{split}0<k_{4}(W_{n})&:=\mathbb{E}(W_{n}^{4})-3(\mathbb{E}(W_{n}^{2}))^{2}\\ &=\frac{48}{n^{2}}\sum_{i,j,k,l=1}^{n}\rho(ih,jh)\rho(jh,kh)\rho(kh,lh)\rho(lh,jh).\end{split}$ (43)
The third and fourth cumulants of $\overline{W}_{n}$ will be similar with $\rho$ replaced by $\rho_{0}$. According to Propositions 6.3 and 6.4 of Biermé et al. [3] and the inequality (5) in Lemma 5.2, we obtain that
$\begin{split}k_{3}(\overline{W}_{n})&\leq\frac{C}{\sqrt{n}}\left(\sum_{|k|<n}\left|\rho_{0}(k)\right|^{\frac{3}{2}}\right)^{2}\leq\frac{C}{\sqrt{n}}\left(\sum_{k=0}^{n-1}(1+k)^{\frac{3}{2}(2H-2)}\right)^{2}\\ &\leq Cn^{(6H-4)\vee 0-\frac{1}{2}}=C\times\left\{\begin{array}[]{ll}\frac{1}{\sqrt{n}},&\quad\text{if }H\in(0,\,\frac{2}{3}],\\ n^{\frac{3}{2}(4H-3)},&\quad\text{if }H\in(\frac{2}{3},\,\frac{3}{4}),\end{array}\right.\end{split}$ (44)
$\begin{split}k_{4}(\overline{W}_{n})&\leq\frac{C}{n}\left(\sum_{|k|<n}\left|\rho_{0}(k)\right|^{\frac{4}{3}}\right)^{3}\leq\frac{C}{n}\left(\sum_{k=0}^{n-1}(1+k)^{\frac{4}{3}(2H-2)}\right)^{3}\\ &\leq Cn^{(8H-5)\vee 0-1}=C\times\left\{\begin{array}[]{ll}\frac{1}{n},&\quad\text{if }H\in(0,\,\frac{5}{8}],\\ n^{2(4H-3)},&\quad\text{if }H\in(\frac{5}{8},\,\frac{3}{4}).\end{array}\right.\end{split}$ (45)
On the other hand, from the identity (18), we rewrite $W_{n}$ as
$\begin{split}W_{n}=\overline{W}_{n}+\frac{1}{\sqrt{n}}\sum_{j=1}^{n}e^{-\theta jh}R_{jh}-\sqrt{n}\left(\mathbb{E}(B_{n})-a\right),\end{split}$ (46)
where $R_{jh}=-2Y_{jh}Y_{0}+e^{-\theta jh}Y_{0}^{2}$, which satisfies
$\begin{split}\sup_{j}\left|\left|R_{j}\right|\right|_{L^{2}(\Omega)}<\infty,\end{split}$ (47)
based on the stationary property of $Y_{t}$ and Cauchy-Schwarz inequality. Combining this result with the fact that $R_{j}$ is a 2-th Wiener chaos, we have
$\begin{split}\sup_{j}\left|\left|\frac{1}{\sqrt{n}}\sum_{j=1}^{n}e^{-\theta jh}R_{jh}\right|\right|_{L^{2}(\Omega)}<\frac{C}{\sqrt{n}},\end{split}$ (48)
where C is independent of $n$. Then, by the identity (46) and Cauchy-Schwarz inequality, Minkowski’s inequality, and hypercontractivity property of Wiener chaos, we obtain that
$\begin{split}&\left|k_{3}({W}_{n})-k_{3}(\overline{W}_{n})\right|=\left|\mathbb{E}\left(W_{n}^{3}-\overline{W}_{n}^{3}\right)\right|\\ &=\left|\mathbb{E}\left[\left(\frac{1}{\sqrt{n}}\sum_{j=1}^{n}e^{-\theta jh}R_{jh}-\sqrt{n}\left(\mathbb{E}(B_{n})-a\right)\right)\left(W_{n}^{2}+W_{n}\overline{W}_{n}+\overline{W}_{n}^{2}\right)\right]\right|\\ &\leq\left[\left|\left|\frac{1}{\sqrt{n}}\sum_{j=1}^{n}e^{-\theta jh}R_{jh}\right|\right|_{L^{2}(\Omega)}+\left|\left|\sqrt{n}\left(\mathbb{E}(B_{n})-a\right)\right|\right|_{L^{2}(\Omega)}\right]\left|\left|W_{n}^{2}+W_{n}\overline{W}_{n}+\overline{W}_{n}^{2}\right|\right|_{L^{2}(\Omega)}\\ &\leq\frac{C}{\sqrt{n}},\end{split}$ (49)
where in the last inequality we also have used the estimation (48) and Propositions 3.1, 3.2. A similar method yields
$\begin{split}&\left|k_{4}({W}_{n})-k_{4}(\overline{W}_{n})\right|=\left|\mathbb{E}\left(W_{n}^{4}\right)-\mathbb{E}\left(\overline{W}_{n}^{4}\right)\right|+3\left|\left(\mathbb{E}W_{n}^{2}\right)^{2}-\left(\mathbb{E}\overline{W}_{n}^{2}\right)^{2}\right|\\ &\leq\left|\left|\frac{1}{\sqrt{n}}\sum_{j=1}^{n}e^{-\theta jh}R_{jh}-\sqrt{n}\left(\mathbb{E}(B_{n})-a\right)\right|\right|_{L^{2}(\Omega)}\cdot\left|\left|W_{n}^{3}+W_{n}^{2}\overline{W}_{n}++W_{n}\overline{W}_{n}^{2}+\overline{W}_{n}^{3}\right|\right|_{L^{2}(\Omega)}\\ &\quad+3\left|\left(\mathbb{E}W_{n}^{2}+\mathbb{E}\overline{W}_{n}^{2}\right)\left(\mathbb{E}W_{n}^{2}-\mathbb{E}\overline{W}_{n}^{2}\right)\right|\\ &\leq\frac{C}{\sqrt{n}}.\end{split}$ (50)
Combining the estimations (44), (45), (49), (50), we can obtain the desired result. ∎
3.2 Berry-Esséen type upper bound for the moment estimator of fOU process
In this section, we concentrate on establishing the Berry-Esséen type upper bound for the moment estimator of the fractional Ornstein-Uhlenbeck process under discrete observations with the fixed step size.
Proof of Theorem 1.1. Recall the definition of $\hat{\theta}_{n}$ and $B_{n}$, we take the random variable $X$, $f(X)$ in the Lemma 2.1 as
$\begin{split}X=B_{n}=\frac{1}{n}\sum_{j=1}^{n}X_{jh}^{2},\quad f(X)=\hat{\theta}_{n}=\left(\frac{1}{{H}\Gamma(2H)}B_{n}\right)^{-\frac{1}{2H}},\end{split}$ (51)
and $a=g(\theta)=H\Gamma(2H)\theta^{-2H}$, $\mathcal{N}=\xi\sim N(0,\sigma_{1}^{2})$, $\eta=\varpi\sim N(0,\sigma_{2}^{2})$. Section 1.3.2.2 of Kubilius et al. [18] shows that $B_{n}\rightarrow{a}$ almost surely, so we have $B_{n}>0$ almost surely. Then, according to Lemma 2.1, there exists a positive constant $C$ independent of $T$ such that for $T$ large enough
$\begin{split}&d_{Kol}\left(\sqrt{n}(\hat{\theta}_{n}-\theta),\mathcal{N}\right)\\ &\leq C\times\left(d_{Kol}(\sqrt{n}(B_{n}-\mathbb{E}[B_{n}]),\varpi)+\sqrt{n}\left|\mathbb{E}[B_{n}]-a\right|+\frac{1}{\sqrt{n}}\right),\end{split}$ (52)
where $\varpi$ is a normal random variable with zero mean and variance $\sigma_{2}^{2}=\sigma^{2}_{B}$ defined as in equation (22) and then $\sigma_{1}^{2}=\frac{\theta^{2}\sigma_{B}^{2}}{4H^{2}a^{2}}$ from (10).
Firstly, we estimate the term $d_{Kol}(\sqrt{n}(B_{n}-\mathbb{E}[B_{n}]),\varpi)$. Denote a sequence of random variables $\varpi_{n}\sim N(0,\sigma_{n}^{2})$ with the variance $\sigma_{n}^{2}=\mathbb{E}(W_{n}^{2})$, where $W_{n}=\sqrt{n}(B_{n}-\mathbb{E}[B_{n}])$ defined in (20). Then, we have that
$\begin{split}d_{Kol}(W_{n},\varpi)\leq d_{Kol}(W_{n},\varpi_{n})+d_{Kol}(\varpi_{n},\varpi),\end{split}$ (53)
by the triangle inequality. The optimal fourth moment theorem of Nourdin and Peccati [22] and the well-known fact that $d_{Kol}(\cdot,\cdot)\leq d_{TV}(\cdot,\cdot)$ imply that
$\begin{split}d_{Kol}(W_{n},\varpi_{n})&\leq d_{TV}(W_{n},\varpi_{n})\leq C\max\left\{k_{3}(W_{n}),k_{4}(W_{n})\right\}\\ &\leq C\times\left\{\begin{array}[]{ll}\frac{1}{\sqrt{n}},&\quad\text{if }H\in(0,\,\frac{2}{3}],\\ n^{\frac{3}{2}(4H-3)},&\quad\text{if }H\in(\frac{2}{3},\,\frac{3}{4}),\end{array}\right.\end{split}$ (54)
where the last inequality is resulted from Proposition 3.3. Using Proposition 3.6.1 of Nourdin and Peccati [21] and Proposition 3.2 yield
$\begin{split}d_{Kol}(\varpi_{n},\varpi)\leq\frac{2}{\sigma_{n}^{2}\vee\sigma_{B}^{2}}\left|\sigma_{n}^{2}-\sigma_{B}^{2}\right|\leq C\times\left\{\begin{array}[]{ll}\frac{1}{n},&\quad\text{if }H\in(0,\,\frac{1}{2}],\\ \frac{1}{n^{3-4H}},&\quad\text{if }H\in(\frac{1}{2},\,\frac{3}{4}).\end{array}\right.\end{split}$ (55)
Combining this result with the inequalities (53), (54) implies that
$\begin{split}(d_{Kol}(\sqrt{n}(B_{n}-\mathbb{E}[B_{n}]),\varpi)\leq C\times\left\{\begin{array}[]{ll}\frac{1}{\sqrt{n}},&\quad\text{if }H\in(0,\,\frac{5}{8}],\\ \frac{1}{n^{3-4H}},&\quad\text{if }H\in(\frac{5}{8},\,\frac{3}{4}).\end{array}\right.\end{split}$ (56)
Secondly, it is straightforward to show for $H\in(0,1)$
$\sqrt{n}\left|\mathbb{E}(B_{n})-a\right|\leq C\times\frac{1}{\sqrt{n}},$ (57)
from Proposition (3.1). Consequently, substituting the inequalities (56), (57) into the estimation (52) yields the desired Berry-Esséen upper bound (3) in Theorem 1.1. $\Box$
4 Proof of Theorem 1.7 and Theorem 1.9
4.1 The second moment and cumulants of the random variable $\widetilde{W}_{n}$
Prior to proving Theorem 1.7, liking the definition of $W_{n}$, we first denote by $\widetilde{W}_{n}$ a second Wiener chaos with respect to the general Gaussian process $G_{t}$ as:
$\widetilde{W}_{n}=\sqrt{n}\left(A_{n}-\mathbb{E}(A_{n})\right)=\frac{1}{\sqrt{n}}\sum_{j=1}^{n}\left(Z_{jh}^{2}-\mathbb{E}\left(Z_{jh}^{2}\right)\right).$ (1)
Establishing the Berry-Esséen upper bound for $\widetilde{W}_{n}$ constitutes a critical step in the proof of Theorem 1.7. The following proposition characterizes the asymptotic behavior of its second moment and convergence rate.
Proposition 4.1.
Let $H\in(0,\frac{1}{2})$ and $\widetilde{W}_{n}$ be defined as in (1). When $n$ large enough, the exist a constant $C$ independent of $n$ such that
$\begin{split}\left|\mathbb{E}(\widetilde{W}_{n}^{2})-\sigma_{B}^{2}\right|\leq C\times\frac{1}{\sqrt{n}},\end{split}$ (2)
where $\sigma_{B}^{2}$ is a series defined in (22).
Proof.
Based on Proposition 3.2, the proof reduces to verifying that
$\begin{split}\left|\mathbb{E}(\widetilde{W}_{n}^{2}-W_{n}^{2})\right|\leq C\times\frac{1}{\sqrt{n}}.\end{split}$ (3)
Denoting the covariance function of $Z_{t}$ as $\tilde{\rho}(t,s)=\mathbb{E}(Z_{t}Z_{s})$, we rewrite the left hand of above inequality as
$\begin{split}\mathbb{E}(\widetilde{W}_{n}^{2}-W_{n}^{2})&=\frac{2}{n}\sum_{j,l=1}^{n}\left[\tilde{\rho}^{2}(jh,lh)-\rho^{2}(jh,lh)\right]\\ &=\frac{2}{n}\sum_{j,l=1}^{n}\left[\left(\tilde{\rho}(jh,lh)-\rho(jh,lh)+\rho(jh,lh)\right)^{2}-\rho^{2}(jh,lh)\right]\\ &=\frac{2}{n}\sum_{j,l=1}^{n}\left(\tilde{\rho}(jh,lh)-\rho(jh,lh)\right)^{2}+\frac{4}{n}\sum_{j,l=1}^{n}\rho(jh,lh)\left(\tilde{\rho}(jh,lh)-\rho(jh,lh)\right)\\ &:=D_{1}+D_{2}.\end{split}$ (4)
Next, we will estimate the upper bound of $D_{1}$ and $D_{2}$, respectively. The estimation (6) in Lemma 5.2, along with the symmetry of $j,l$ and the change variable $k=l-j$ imply that
$\begin{split}0\leq D_{1}&\leq\frac{C}{n}\sum_{j,l=1}^{n}\left[\left(1+(j\wedge l)\right)^{2(H-1)}\wedge\left(1+\left|j-l\right|\right)^{H-1}\right]^{2}\\ &\leq\frac{C}{n}\sum_{1\leq j\leq l\leq n}\left[\left(1+j\right)^{2(H-1)}\wedge\left(1+(l-j)\right)^{H-1}\right]^{2}\\ &\leq\frac{C}{n}\sum_{1\leq j\leq l\leq n}\left(1+j\right)^{2(H-1)}\cdot\left(1+(l-j)\right)^{H-1}\\ &\leq\frac{C}{n}\sum_{j=1}^{n}\left(1+j\right)^{2(H-1)}\sum_{k=0}^{n-1}\left(1+k\right)^{H-1}\\ &\leq\frac{C}{n}n^{(2H-1)\vee 0}n^{H}=Cn^{H-1}\leq C\frac{1}{\sqrt{n}},\end{split}$ (5)
where the last two inequalities are from condition $H\in(0,\frac{1}{2})$. Using the triangle inequality, estimations (5), (6) in Lemma 5.2 and the symmetry of $j,l$ yield that
$\begin{split}\left|D_{2}\right|&\leq\frac{C}{n}\sum_{j,l=1}^{n}\left(1+\left|j-l\right|\right)^{2(H-1)}\cdot\left[\left(1+(j\wedge l)\right)^{2(H-1)}\wedge\left(1+\left|j-l\right|\right)^{H-1}\right]\\ &\leq\frac{C}{n}\sum_{1\leq j\leq l\leq n}\left(1+(l-j)\right)^{2(H-1)}\cdot\left(1+j\right)^{2(H-1)}\\ &\leq\frac{C}{n}\left(\sum_{j=1}^{n}\left(1+j\right)^{2(H-1)}\right)^{2}\leq\frac{C}{n}n^{(4H-2)\vee 0}=\frac{C}{n}.\end{split}$ (6)
Consequently, substituting inequality (5), (6) into equation (4) obtains the desired result (3). In summary, this completes the proof. ∎
Next, we focus on estimating the third and fourth cumulants of random variable $\widetilde{W}_{n}$.
Proposition 4.2.
Let $H\in(0,\frac{1}{2})$ and $\widetilde{W}_{n}$ be defined as in (1). Denote the third cumulants of random variable $\widetilde{W}_{n}$ by
$\begin{split}k_{3}(\widetilde{W}_{n}):=\mathbb{E}(\widetilde{W}_{n}^{3})=\frac{8}{n^{3/2}}\sum_{j,k,l=1}^{n}\tilde{\rho}(jh,kh)\tilde{\rho}(kh,lh)\tilde{\rho}(lh,jh).\end{split}$ (7)
When $n$ large enough, the exist a constant $C$ independent of $n$ such that
$\begin{split}\left|k_{3}(\widetilde{W}_{n})\right|\leq C\times\frac{1}{\sqrt{n}}.\end{split}$ (8)
Proof.
According to the estimations (44) and (49) in Proposition 3.3, we only need to show
$\begin{split}\left|k_{3}(\widetilde{W}_{n})-k_{3}(W_{n})\right|\leq C\times\frac{1}{\sqrt{n}},\end{split}$ (9)
The above inequality is equivalent to
$\begin{split}I:=\left|\sum_{j,k,l=1}^{n}\Big{[}\tilde{\rho}(jh,kh)\tilde{\rho}(kh,lh)\tilde{\rho}(lh,jh)-\rho(jh,kh)\rho(kh,lh)\rho(lh,jh)\Big{]}\right|\leq Cn.\end{split}$ (10)
For the sake of simplicity, we denote $x=\tilde{\rho}(jh,kh)-\rho(jh,kh),y=\tilde{\rho}(kh,lh)-\rho(kh,lh),z=\tilde{\rho}(lh,jh)-\rho(lh,jh)$, then $I$ is decomposed into the following seven summations,
$\begin{split}I&=\Bigg{|}\sum_{j,k,l=1}^{n}\Big{[}x\rho(kh,lh)\rho(lh,jh)+\rho(jh,kh)y\rho(lh,jh)+\rho(jh,kh)\rho(kh,lh)z\\ &\qquad\qquad\quad+xy\rho(lh,jh)+x\rho(kh,lh)z+\rho(jh,kh)yz+xyz\Big{]}\Bigg{|}\\ &:=\left|\sum_{i=1}^{7}I_{i}\right|.\end{split}$ (11)
Next, we estimate the upper bound for each of $I_{i},i=1,\cdots,7$. The key point is to select the scaling approach of $x,y,z$ based on the different symmetries of $i,j,k$ in every $I_{i}$. The estimations (5), (6) in Lemma 5.2 and the symmetry of $j,k$ imply that
$\begin{split}\left|I_{1}\right|&\leq C\sum_{j,k,l=1}^{n}\left|x\rho(kh,lh)\rho(lh,jh)\right|\\ &\leq C\sum_{j,k,l=1}^{n}\left(1+(j\wedge k)\right)^{2(H-1)}\cdot\left(1+\left|k-l\right|\right)^{2(H-1)}\cdot\left(1+\left|l-j\right|\right)^{2(H-1)}\\ &\leq C\sum_{1\leq j\leq k\leq n,1\leq l\leq n}\left(1+j\right)^{2(H-1)}\cdot\left(1+\left|k-l\right|\right)^{2(H-1)}\cdot\left(1+\left|l-j\right|\right)^{2(H-1)}\\ &\leq C,\end{split}$ (12)
where in the last inequality we use Lemma 5.3 with the condition $H\in(0,\frac{1}{2})$.
With the similar way, we also have
$\begin{split}\left|I_{2}\right|\leq C\sum_{j,k,l=1}^{n}\left(1+\left|j-k\right|\right)^{2(H-1)}\cdot\left(1+(k\wedge l)\right)^{2(H-1)}\cdot\left(1+\left|l-j\right|\right)^{2(H-1)}\leq C,\end{split}$ (13)
$\begin{split}\left|I_{3}\right|\leq C\sum_{j,k,l=1}^{n}\left(1+\left|j-k\right|\right)^{2(H-1)}\cdot\left(1+\left|k-l\right|\right)^{2(H-1)}\cdot\left(1+(l\wedge j)\right)^{2(H-1)}\leq C.\end{split}$ (14)
According to Lemma 5.2 and the symmetry of $l,j$, we can derive that
$\begin{split}\left|I_{4}\right|&\leq C\sum_{j,k,l=1}^{n}\left|xy\rho(lh,jh)\right|\\ &\leq C\sum_{j,k,l=1}^{n}\left(1+\left|j-k\right|\right)^{H-1}\cdot\left(1+\left|k-l\right|\right)^{H-1}\cdot\left(1+\left|l-j\right|\right)^{2(H-1)}\\ &\leq C\sum_{1\leq l\leq j\leq n,1\leq k\leq n}\left(1+\left|j-k\right|\right)^{H-1}\cdot\left(1+\left|k-l\right|\right)^{H-1}\cdot\left(1+(j-l)\right)^{2(H-1)}\\ &\leq Cn^{2H}\leq Cn,\end{split}$ (15)
where in the last inequality we use Lemma 5.3 with the condition $H\in(0,\frac{1}{2})$.
With the similar way, we also have
$\begin{split}\left|I_{5}\right|\leq C\sum_{j,k,l=1}^{n}\left(1+\left|j-k\right|\right)^{H-1}\cdot\left(1+\left|k-l\right|\right)^{2(H-1)}\cdot\left(1+\left|l-j\right|\right)^{H-1}&\leq Cn^{2H}\leq Cn,\end{split}$ (16)
$\begin{split}\left|I_{6}\right|\leq C\sum_{j,k,l=1}^{n}\left(1+\left|j-k\right|\right)^{2(H-1)}\cdot\left(1+\left|k-l\right|\right)^{H-1}\cdot\left(1+\left|l-j\right|\right)^{H-1}&\leq Cn^{2H}\leq Cn.\end{split}$ (17)
For the last term $I_{7}$, we use Lemma 5.2 and the symmetry of $j,k,l$ obtain that
$\begin{split}\left|I_{7}\right|&\leq C\sum_{j,k,l=1}^{n}\left|xyz\right|\leq C\sum_{j,k,l=1}^{n}\left(1+(j\wedge k)\right)^{2(H-1)}\cdot\left(1+(k\wedge l)\right)^{2(H-1)}\cdot\left(1+\left|l-j\right|\right)^{H-1}\\ &\leq C\sum_{1\leq j\leq k\leq l\leq n}\left(1+j\right)^{2(H-1)}\cdot\left(1+k\right)^{2(H-1)}\cdot\left(1+(l-j)\right)^{H-1}\leq Cn^{H}\leq Cn,\end{split}$ (18)
where the last inequality is from Lemma 5.3 with the condition $H\in(0,\frac{1}{2})$. In conclusion, we get
$\begin{split}0\leq I\leq Cn.\end{split}$ (19)
This complete the proof. ∎
Proposition 4.3.
Let $H\in(0,\frac{1}{2})$ and $\widetilde{W}_{n}$ be defined as in (1). Denote the forth cumulants of random variable $\widetilde{W}_{n}$ by
$\begin{split}0<k_{4}(\widetilde{W}_{n})&:=\mathbb{E}(\widetilde{W}_{n}^{4})-3(\mathbb{E}(\widetilde{W}_{n}^{2}))^{2}\\ &=\frac{48}{n^{2}}\sum_{i,j,k,l=1}^{n}\tilde{\rho}(ih,jh)\tilde{\rho}(jh,kh)\tilde{\rho}(kh,lh)\tilde{\rho}(lh,jh).\end{split}$ (20)
When $n$ is large enough, the exist a constant $C$ independent of $n$ such that
$\begin{split}k_{4}(\widetilde{W}_{n})\leq C\times\frac{1}{\sqrt{n}}.\end{split}$ (21)
Proof.
From the previous estimation concerning $k_{4}(\overline{W}_{n})$ and $\left|k_{4}({W}_{n})-k_{4}(\overline{W}_{n})\right|$ in Proposition 3.3, it is sufficient to prove that
$\begin{split}\left|k_{4}(\widetilde{W}_{n})-k_{4}(W_{n})\right|\leq C\times\frac{1}{\sqrt{n}},\end{split}$ (22)
which is equivalent to showing
$\begin{split}J&:=\left|\sum_{j,k,l,m=1}^{n}\Big{[}\tilde{\rho}(jh,kh)\tilde{\rho}(kh,lh)\tilde{\rho}(lh,mh)\tilde{\rho}(mh,jh)-\rho(jh,kh)\rho(kh,lh)\rho(lh,mh)\rho(mh,jh)\Big{]}\right|\\ &\leq Cn^{\frac{3}{2}}.\end{split}$ (23)
Denoting by $x,y$ the same symbol as in Proposition 4.2 and $z=\tilde{\rho}(lh,mh)-\rho(lh,mh),w=\tilde{\rho}(mh,jh)-\rho(mh,jh)$, which implies that $J$ can be decomposed into fifteen summations:
$\begin{split}J&=\Bigg{|}\sum_{j,k,l,m=1}^{n}\Big{[}x\rho(kh,lh)\rho(lh,mh)\rho(mh,jh)+\rho(jh,kh)y\rho(lh,mh)\rho(mh,jh)\\ &\qquad\qquad\quad+\rho(jh,kh)\rho(kh,lh)z\rho(mh,jh)+\rho(jh,kh)\rho(kh,lh)\rho(lh,mh)w\\ &\qquad\qquad\quad+xy\rho(lh,mh)\rho(mh,jh)+x\rho(kh,lh)z\rho(mh,jh)+x\rho(kh,lh)\rho(lh,mh)w\\ &\qquad\qquad\quad+\rho(jh,kh)yz\rho(mh,jh)+\rho(jh,kh)y\rho(lh,mh)w+\rho(jh,kh)\rho(kh,lh)zw\\ &\qquad\qquad\quad+xyz\rho(mh,jh)+xy\rho(lh,mh)w+x\rho(kh,lh)zw+\rho(jh,kh)yzw+xyzw\Big{]}\Bigg{|}\\ &:=\Bigg{|}\sum_{i=1}^{15}J_{i}\Bigg{|}.\end{split}$ (24)
We divide the fifteen summations into five groups and discuss them separately. The idea is to analyze each term $J_{i}$ by the different symmetry of the sum index in $J_{i}$.
Group 1. Lemma 5.2 and the symmetry of $j,k$ imply that
$\begin{split}&\left|J_{1}\right|=\left|\sum_{j,k,l,m=1}^{n}x\rho(kh,lh)\rho(lh,mh)\rho(mh,jh)\right|\\ &\leq C\sum_{j,k,l,m=1}^{n}\left(1+(j\wedge k)\right)^{2(H-1)}\cdot\left(1+\left|k-l\right|\right)^{2(H-1)}\cdot\left(1+\left|l-m\right|\right)^{2(H-1)}\cdot\left(1+\left|m-j\right|\right)^{2(H-1)}\\ &\leq C\sum_{1\leq j\leq k\leq n,1\leq l,m\leq n}\left(1+j\right)^{2(H-1)}\cdot\left(1+\left|k-l\right|\right)^{2(H-1)}\cdot\left(1+\left|l-m\right|\right)^{2(H-1)}\cdot\left(1+\left|m-j\right|\right)^{2(H-1)}\\ &\leq C,\end{split}$ (25)
where in the last inequality we use Lemma 5.3 with the condition $H\in(0,\frac{1}{2})$. Similarly, using symmetry and analogous arguments, we obtain:
$\begin{split}\left|J_{i}\right|\leq C,\quad j=2,3,4.\end{split}$ (26)
Group 2. For the term $J_{5}$, noticing the symmetry of $j,l$ and utilizing Lemma 5.2, we have
$\left|J_{5}\right|=\left|\sum_{j,k,l,m=1}^{n}xy\rho(lh,mh)\rho(mh,jh)\right|$
$\leq C\sum_{j,k,l,m=1}^{n}\left(1+(j\wedge k)\right)^{2(H-1)}\cdot\left(1+\left|k-l\right|\right)^{H-1}\cdot\left(1+\left|l-m\right|\right)^{2(H-1)}\cdot\left(1+\left|m-j\right|\right)^{2(H-1)}$
$\leq C\Bigg{[}\sum_{1\leq k\leq j\leq l\leq n,1\leq m\leq n}\left(1+k\right)^{2(H-1)}\cdot\left(1+(l-k)\right)^{H-1}\cdot\left(1+\left|l-m\right|\right)^{2(H-1)}\cdot\left(1+\left|m-j\right|\right)^{2(H-1)}$
$+\sum_{1\leq j\leq k\leq l\leq n,1\leq m\leq n}\left(1+j\right)^{2(H-1)}\cdot\left(1+(l-k)\right)^{H-1}\cdot\left(1+\left|l-m\right|\right)^{2(H-1)}\cdot\left(1+\left|m-j\right|\right)^{2(H-1)}$
$+\sum_{1\leq j\leq l\leq k\leq n,1\leq m\leq n}\left(1+j\right)^{2(H-1)}\cdot\left(1+(k-l)\right)^{H-1}\cdot\left(1+\left|l-m\right|\right)^{2(H-1)}\cdot\left(1+\left|m-j\right|\right)^{2(H-1)}\Bigg{]}$
$\leq Cn^{H},$ (27)
where the last inequality is caused by Lemma 5.3 with the condition $H\in(0,\frac{1}{2})$. Notice that $J_{i},j=7,8,10$ share the similar symmetry with $J_{5}$, which implies that
$\begin{split}\left|J_{i}\right|\leq Cn^{H},\quad j=7,8,10.\end{split}$ (28)
Group 3. We estimate the term $J_{6}$ by the symmetry of $(j,k)$ and $(l,m)$ with Lemma 5.2,
$\begin{split}&\left|J_{6}\right|=\left|\sum_{j,k,l,m=1}^{n}x\rho(kh,lh)z\rho(mh,jh)\right|\\ &\leq C\sum_{j,k,l,m=1}^{n}\left(1+(j\wedge k)\right)^{2(H-1)}\cdot\left(1+\left|k-l\right|\right)^{2(H-1)}\cdot\left(1+\left|l-m\right|\right)^{H-1}\cdot\left(1+\left|m-j\right|\right)^{2(H-1)}\\ &\leq C\Bigg{[}\sum_{1\leq j\leq k\leq n,1\leq l\leq m\leq n}\left(1+j\right)^{2(H-1)}\cdot\left(1+\left|k-l\right|\right)^{2(H-1)}\cdot\left(1+(m-l)\right)^{H-1}\cdot\left(1+\left|m-j\right|\right)^{2(H-1)}\\ &+\sum_{1\leq j\leq k\leq n,1\leq m\leq l\leq n}\left(1+j\right)^{2(H-1)}\cdot\left(1+\left|k-l\right|\right)^{2(H-1)}\cdot\left(1+(l-m)\right)^{H-1}\cdot\left(1+\left|m-j\right|\right)^{2(H-1)}\Bigg{]}\\ &\leq Cn^{H},\end{split}$ (29)
where the last inequality is from Lemma 5.3 with the condition $H\in(0,\frac{1}{2})$. At the same time, $J_{9}$ has the symmetry of $(k,l)$ and $(m,j)$. By a similar way we obtain
$\begin{split}\left|J_{9}\right|\leq Cn^{H}.\end{split}$ (30)
Group 4. Applying the symmetry of $m,j$ and Lemma 5.2 to $J_{11}$:
$\left|J_{11}\right|=\left|\sum_{j,k,l,m=1}^{n}xyz\rho(mh,jh)\right|$
$\leq C\sum_{j,k,l,m=1}^{n}\left(1+(j\wedge k)\right)^{2(H-1)}\cdot\left(1+\left|k-l\right|\right)^{H-1}\cdot\left(1+\left|l-m\right|\right)^{H-1}\cdot\left(1+\left|m-j\right|\right)^{2(H-1)}$
$\leq C\Bigg{[}\sum_{1\leq m\leq k\leq j\leq n,1\leq l\leq n}\left(1+k\right)^{2(H-1)}\cdot\left(1+\left|k-l\right|\right)^{H-1}\cdot\left(1+\left|l-m\right|\right)^{H-1}\cdot\left(1+(j-m)\right)^{2(H-1)}$
$+\sum_{1\leq m\leq j\leq k\leq n,1\leq l\leq n}\left(1+j\right)^{2(H-1)}\cdot\left(1+\left|k-l\right|\right)^{H-1}\cdot\left(1+\left|l-m\right|\right)^{H-1}\cdot\left(1+(j-m)\right)^{2(H-1)}$
$+\sum_{1\leq k\leq m\leq j\leq n,1\leq l\leq n}\left(1+k\right)^{2(H-1)}\cdot\left(1+\left|k-l\right|\right)^{H-1}\cdot\left(1+\left|l-m\right|\right)^{H-1}\cdot\left(1+(j-m)\right)^{2(H-1)}\Bigg{]}$
$\leq Cn^{2H},$ (31)
where the last inequality is due to Lemma 5.3 with the condition $H\in(0,\frac{1}{2})$. Similarly, we have
$\begin{split}\left|J_{i}\right|\leq Cn^{2H},\quad i=12,13,14.\end{split}$ (32)
Group 5. We using the symmetry of $j,k,l,m$ and Lemma 5.2 to get that
$\begin{split}&\left|J_{15}\right|=\left|\sum_{j,k,l,m=1}^{n}xyzw\right|\\ &\leq C\sum_{j,k,l,m=1}^{n}\left(1+(j\wedge k)\right)^{2(H-1)}\cdot\left(1+(k\wedge l)\right)^{2(H-1)}\cdot\left(1+(l\wedge m)\right)^{2(H-1)}\cdot\left(1+\left|m-j\right|\right)^{H-1}\\ &\leq C\sum_{1\leq j\leq k\leq l\leq m\leq n}\left(1+j\right)^{2(H-1)}\cdot\left(1+k\right)^{2(H-1)}\cdot\left(1+l\right)^{2(H-1)}\cdot\left(1+(m-j)\right)^{H-1}\\ &\leq Cn^{H},\end{split}$ (33)
where the last inequality is from Lemma 5.3 with the condition $H\in(0,\frac{1}{2})$. In conclusion, we obtain
$\begin{split}J\leq\sum_{i=1}^{15}\left|J_{i}\right|\leq Cn^{2H}\leq Cn^{\frac{3}{2}}.\end{split}$ (34)
This completes the proof. ∎
4.2 Proofs of main Theorems
Proof of Theorem 1.7. Following the proof methodology of Theorem 1.1, we take the random variable $X$ as the second moment of sample about Ornstein-Uhlenbeck model $\{Z_{t}:t\geq 0\}$ defined as in (4) with the discrete form:
$X=A_{n}:=\frac{1}{n}\sum_{j=1}^{n}Z_{jh}^{2}.$ (35)
Lemma 2.1 is also a key tool for proving Berry-Esséen upper bound of $\hat{\theta}_{n}$ defined in (7), which implies that there exists a positive constant $C$ independent of $n$ such that for $n$ large enough
$\begin{split}&d_{Kol}\left(\sqrt{n}(\hat{\theta}_{n}-\theta),\mathcal{N}\right)\\ &\leq C\times\left(d_{Kol}(\sqrt{n}(A_{n}-\mathbb{E}[A_{n}]),\varpi)+\sqrt{n}\left|\mathbb{E}[A_{n}]-a\right|+\frac{1}{\sqrt{n}}\right),\end{split}$ (36)
where $a=g(\theta)=H\Gamma(2H)\theta^{-2H}$ and $\mathcal{N},\,\varpi$ are the same as in estimation (52). Throughout this proof, we assume $H\in(0,\frac{1}{2})$.
To estimate $\sqrt{n}\left|\mathbb{E}[A_{n}]-a\right|$, we first note from (57) that it suffices to prove that $\sqrt{n}\left|\mathbb{E}(B_{n}-A_{n})\right|\leq C\times\frac{1}{\sqrt{n}}$. In fact, the inequality (7) in Lemma 5.2 implies that
$\begin{split}\sqrt{n}\left|\mathbb{E}(B_{n}-A_{n})\right|&\leq\frac{1}{\sqrt{n}}\sum_{j=1}^{n}\left|\mathbb{E}\left(Z_{jh}^{2}-X_{jh}^{2}\right)\right|\leq\frac{C}{\sqrt{n}}\sum_{j=1}^{n}\left(1\wedge(jh)^{2(H-1)}\right)\\ &\leq\frac{C}{\sqrt{n}}\sum_{j=1}^{n}\left(1+jh\right)^{2(H-1)}\leq C\times n^{(2H-1)\vee 0-\frac{1}{2}}=C\times\frac{1}{\sqrt{n}}.\end{split}$ (37)
Next, using arguments analogous to those in the estimation of (56) combined with Propositions 4.1, 4.2, 4.3, we derive that for $H\in(0,\frac{1}{2})$,
$\begin{split}d_{Kol}(\sqrt{n}(A_{n}-\mathbb{E}[A_{n}]),\varpi)\leq C\times\frac{1}{\sqrt{n}}.\end{split}$ (38)
Substituting the estimations (37), (38) into (36) implies the final Berry-Esséen upper bound
$d_{Kol}\left(\sqrt{n}(\hat{\theta}_{n}-\theta),\mathcal{N}\right)\leq C_{\theta,H,h}\times\frac{1}{\sqrt{n}}.$ (39)
$\Box$
Proof of Theorem 1.9. The distinction between Theorem 1.7 and Theorem 1.9 lies in improving Hypothesis 1.6 with $H\in(0,\frac{1}{2})$ to Hypothesis 1.8 with $H\in(\frac{1}{2},\frac{3}{4})$, thereby accommodating broader Gaussian processes such as the bi-fractional Brownian motion and the sub-bifractional Brownian motion. Notice that the estimations in Lemma 5.2 play a key rule in the proof of Theorem 1.7. To establish Theorem 1.9, it suffices to build up a new comparison about covariance functions ${\rho}(t,s)$ and $\tilde{\rho}(t,s)$, which is presented in the following Proposition.
Proposition 4.4.
Let ${\rho}(t,s)=\mathbb{E}(X_{t}X_{s})$, $\tilde{\rho}(t,s)=\mathbb{E}(Z_{t}Z_{s})$ be the covariance function of the Ornstein-Uhlenbeck processes $X_{t}$ and $Z_{t}$ driven by fBm $B_{t}^{H}$ and $G_{t}$ satisfying Hypothese 1.8 with $H\in(\frac{1}{2},1)$. Then there exists a constant $C\geq 0$ independent of $T$ such that for any $0\leq s\leq t\leq T$,
$\left|\tilde{\rho}(t,s)-\rho(t,s)\right|\leq C\left(1\wedge s^{2(H-1)}\wedge(t-s)^{2(H-1)}\right).$ (40)
Moreover, the difference of variance of $X_{t}$ and $Z_{t}$ satisfies
$\left|\mathbb{E}[Z_{t}^{2}]-\mathbb{E}[X_{t}^{2}]\right|\leq C\left(1\wedge t^{2(H-1)}\right).$ (41)
Proof.
For any $0\leq s\leq t\leq T$, according to the relationship (6) between the inner products of two functions in the Hilbert spaces $\mathfrak{H}$ and $\mathfrak{H}_{1}$, we have
$\left|\mathbb{E}[Z_{t}^{2}]-\mathbb{E}[X_{t}^{2}]\right|\leq\int_{0}^{t}e^{-\theta(t-u)}\mathrm{d}u\int_{0}^{t}e^{-\theta(t-v)}\left|\frac{\partial^{2}R(u,v)}{\partial u\partial v}-\frac{\partial^{2}R^{B}(u,v)}{\partial u\partial v}\right|\mathrm{d}v.$ (42)
At the same time, Hypothesis 1.8 with restriction $H\in(\frac{1}{2},1)$ implies that
$\left|\frac{\partial}{\partial s}\left(\frac{\partial R(s,t)}{\partial t}-\frac{\partial R^{B}(s,t)}{\partial t}\right)\right|\leq C_{1}(t+s)^{2H-2}+C_{2}(s^{2H^{\prime}}+t^{2H^{\prime}})^{K-2}(st)^{2H^{\prime}-1},$ (43)
where $H^{\prime}\in(\frac{1}{2},1),\,K\in(0,2)$ and $H:=H^{\prime}K\in(\frac{1}{2},1)$. Then, by the basic inequality $a+b\geq 2\sqrt{ab}$ and Lemma 5.1, we obtain that
$\begin{split}\left|\mathbb{E}[Z_{t}^{2}]-\mathbb{E}[X_{t}^{2}]\right|&\leq C_{1}\int_{0}^{t}e^{-\theta(t-u)}\mathrm{d}u\int_{0}^{t}e^{-\theta(t-v)}(u+v)^{2H-2}\mathrm{d}v\\ &+C_{2}\int_{0}^{t}e^{-\theta(t-u)}\mathrm{d}u\int_{0}^{t}e^{-\theta(t-v)}(u^{2H^{\prime}}+v^{2H^{\prime}})^{K-2}(uv)^{2H^{\prime}-1}\mathrm{d}v\\ &\leq C_{1}\left[\int_{0}^{t}e^{-\theta(t-u)}u^{H-1}\mathrm{d}u\right]^{2}+C_{2}\left[\int_{0}^{t}e^{-\theta(t-u)}u^{H^{\prime}k-1}\mathrm{d}u\right]^{2}\\ &\leq C\left(1\wedge t^{2(H-1)}\right),\end{split}$ (44)
which proves the inequality (42). Furthermore, combining this result with the equations (3.37) and (3.41) of [8] yield
$\begin{split}&\left|\tilde{\rho}(t,s)-\rho(t,s)\right|\\ &\leq Ce^{-\theta(t-s)}\left|\mathbb{E}[Z_{s}^{2}]-\mathbb{E}[X_{s}^{2}]\right|+\int_{0}^{s}e^{-\theta(s-u)}\mathrm{d}u\int_{s}^{t}e^{-\theta(t-v)}\left|\frac{\partial^{2}R(u,v)}{\partial u\partial v}-\frac{\partial^{2}R^{B}(u,v)}{\partial u\partial v}\right|\mathrm{d}v\\ &\leq C\left(1\wedge s^{2(H-1)}\wedge(t-s)^{2(H-1)}\right)+\int_{0}^{s}e^{-\theta(s-u)}\mathrm{d}u\int_{s}^{t}e^{-\theta(t-v)}\left|\frac{\partial^{2}R(u,v)}{\partial u\partial v}-\frac{\partial^{2}R^{B}(u,v)}{\partial u\partial v}\right|\mathrm{d}v\end{split}$ (45)
where the last inequality is from the fact $e^{-\theta(t-s)}\leq C\left(1\wedge(t-s)^{2(H-1)}\right)$.
Next, we define a double integral as
$\begin{split}{\rm II}&:=\int_{0}^{s}e^{-\theta(s-u)}\mathrm{d}u\int_{s}^{t}e^{-\theta(t-v)}\left|\frac{\partial^{2}R(u,v)}{\partial u\partial v}-\frac{\partial^{2}R^{B}(u,v)}{\partial u\partial v}\right|\mathrm{d}v\\ &\leq C_{1}\int_{0}^{s}e^{-\theta(s-u)}\mathrm{d}u\int_{s}^{t}e^{-\theta(t-v)}(u+v)^{2H-2}\mathrm{d}v\\ &+C_{2}\int_{0}^{s}e^{-\theta(s-u)}\mathrm{d}u\int_{s}^{t}e^{-\theta(t-v)}(u^{2H^{\prime}}+v^{2H^{\prime}})^{K-2}(uv)^{2H^{\prime}-1}\mathrm{d}v\\ &:={\rm II}_{1}+{\rm II}_{2},\end{split}$ (46)
under Hypothesis 1.8. Let’s estimate ${\rm II}$ in two parts based on the inequality (43).
Part 1. Making the change of variable $x=v-s$ implies that
$\begin{split}{\rm II}_{1}\leq C\int_{0}^{s}e^{-\theta(s-u)}\mathrm{d}u\int_{0}^{t-s}e^{-\theta((t-s)-x)}x^{2H-2}\mathrm{d}x\leq C(t-s)^{2H-2}\end{split}$ (47)
where in the last inequality we use Lemma 6 with $H\in(\frac{1}{2},1)$.
Part 2. Suppose that $H^{\prime}\in(\frac{1}{2},1),\,K\in(0,2)$ and $H:=H^{\prime}K\in(\frac{1}{2},1)$. We make the change of variable $x=v-s$ and use the L’Hôpital’s rule to get that
$\begin{split}&\lim_{y\rightarrow\infty}\frac{{\rm II}_{2}}{y^{2(H^{\prime}K-1)}}\\ &\leq C\lim_{y\rightarrow\infty}\frac{\int_{0}^{s}e^{-\theta(s-u)}u^{2H^{\prime}-1}\mathrm{d}u\int_{0}^{y}e^{\theta x}(u^{2H^{\prime}}+(x+s)^{2H^{\prime}})^{K-2}(x+s)^{2H^{\prime}-1}\mathrm{d}x}{y^{2(H^{\prime}K-1)}e^{\theta y}}\\ &=C\lim_{y\rightarrow\infty}\frac{\int_{0}^{s}e^{-\theta(s-u)}u^{2H^{\prime}-1}\mathrm{d}u(u^{2H^{\prime}}+(y+s)^{2H^{\prime}})^{K-2}(y+s)^{2H^{\prime}-1}}{\theta y^{2(H^{\prime}K-1)}+2(H^{\prime}K-1)y^{2(H^{\prime}K-1)-1}}\\ &=C\lim_{y\rightarrow\infty}\int_{0}^{s}e^{-\theta(s-u)}(\frac{u}{y})^{2H^{\prime}-1}\left[(\frac{u}{y})^{2H^{\prime}}+(1+\frac{s}{y})^{2H^{\prime}}\right]^{K-2}\mathrm{d}u\frac{(1+\frac{s}{y})^{2H^{\prime}-1}}{\theta+2(H^{\prime}K-1)y^{-1}},\end{split}$ (48)
where the last equality is from $2(H^{\prime}K-1)=2H^{\prime}(K-2)+2(2H^{\prime}-1)$. Notice that
$\begin{split}\lim_{y\rightarrow\infty}\frac{(1+\frac{s}{y})^{2H^{\prime}-1}}{\theta+2(H^{\prime}K-1)y^{-1}}=\frac{1}{\theta}.\end{split}$ (49)
And then, because $u\in(0,s)$, $s$ is fixed, choosing $y>s$, we have
$(\frac{u}{y})^{2H^{\prime}-1}\left[(\frac{u}{y})^{2H^{\prime}}+(1+\frac{s}{y})^{2H^{\prime}}\right]^{K-2}\leq 1.$
The Lebesgue dominated convergence theorem yields
$\begin{split}\lim_{y\rightarrow\infty}\int_{0}^{s}e^{-\theta(s-u)}(\frac{u}{y})^{2H^{\prime}-1}\left[(\frac{u}{y})^{2H^{\prime}}+(1+\frac{s}{y})^{2H^{\prime}}\right]^{K-2}\mathrm{d}u=0,\end{split}$ (50)
with the condition $H^{\prime}\in(\frac{1}{2},1)$. Consequently, we have
$\begin{split}\lim_{y\rightarrow\infty}\frac{{\rm II}_{2}}{y^{2(H^{\prime}K-1)}}<+\infty,\end{split}$ (51)
which means that
$\begin{split}{\rm II}_{2}\leq C(t-s)^{2(H^{\prime}K-1)}=C(t-s)^{2(H-1)}.\end{split}$ (52)
In conclusion, we obtain that
$\left|\tilde{\rho}(t,s)-\rho(t,s)\right|\leq C\left(1\wedge s^{2(H-1)}\wedge(t-s)^{2(H-1)}\right).$ (53)
Based on the results of Proposition 4.4, the conclusion of Theorem 1.9 can be readily verified via an approach parallel to that of Theorem 1.7. This completes the proof. $\Box$
5 Appendix
We have been used the following technical inequalities repeatedly throughout the paper, which is cited from Chen et al. [7, 4, 8].
Lemma 5.1.
Assume $\beta>-1$, $\theta>0$ and two functions with form
$\begin{split}A_{1}(t)=\int_{0}^{t}e^{-\theta x}x^{\beta}\mathrm{d}x,\qquad A_{2}(t)=\int_{0}^{t}e^{-\theta(t-x)}x^{\beta}\mathrm{d}x,\end{split}$ (1)
then there exist a positive constant $C$ such that for any $s\in\left[0,\infty\right)$,
$\begin{split}A_{1}(t)\leq C(t^{\beta+1}{\bf 1}_{[0,1]}(t)+{\bf 1}_{(1,\infty)}(t))\leq C(1\wedge t^{\beta+1}),\end{split}$ (2)
$\begin{split}A_{2}(t)\leq C(t^{\beta+1}{\bf 1}_{[0,1]}(t)+t^{\beta}{\bf 1}_{(1,\infty)}(t))\leq C(t^{\beta}\wedge t^{\beta+1}).\end{split}$ (3)
In particular, if $\beta\in(-1,0)$, then there exist a positive constant $C$ such that for any $s\in\left[0,\infty\right)$,
$\begin{split}A_{2}(t)\leq C(1\wedge t^{\beta}).\end{split}$ (4)
Lemma 5.2.
Denote
${\rho}(t,s)=\mathbb{E}(X_{t}X_{s}),\quad\tilde{\rho}(t,s)=\mathbb{E}(Z_{t}Z_{s})$
by the covariance function of the Ornstein-Uhlenbeck processes $X_{t}$ and $Z_{t}$ driven by fBm $B_{t}^{H}$ and $G_{t}$ satisfying Hypothese 1.6. Then there exists a positive constant $C$ independent of $T$ such that for any $0\leq s\leq t\leq T$,
$\left|\rho(t,s)\right|\leq C\left(1\wedge(t-s)^{2(H-1)}\right)\leq C\left(1+(t-s)\right)^{2(H-1)},$ (5)
$\left|\tilde{\rho}(t,s)-\rho(t,s)\right|\leq C\left(1\wedge s^{2(H-1)}\wedge(t-s)^{H-1}\right).$ (6)
Moreover, the difference of variance of $X_{t}$ and $Z_{t}$ satisfies
$\left|\mathbb{E}[Z_{t}^{2}]-\mathbb{E}[X_{t}^{2}]\right|\leq C\left(1\wedge s^{2(H-1)}\right)\leq C\left(1+s\right)^{2(H-1)}.$ (7)
Lemma 5.3.
If $r\in\mathbb{N}:=\{1,2,\cdots\}$ is large enough and $v_{1},\cdots,v_{l}$ are positive, then there exists a positive constant $C$ depending on $v_{1},\cdots,v_{l}$ such that
$\sum_{r_{i}\in\mathbb{N},\sum_{i=1}^{l}r_{i}<r}r_{1}^{v_{1}-1}r_{2}^{v_{2}-1}\cdots r_{l}^{v_{l}-1}\leq C\times r^{\sum_{i=1}^{l}v_{i}}.$ (8)
At the same time, if $r\in\mathbb{N}:=\{1,2,\cdots\}$ is large enough and $v_{1},\cdots,v_{l}$ are negative, then there exists a positive constant $C$ depending on $v_{1},\cdots,v_{l}$ such that
$\sum_{r_{i}\in\mathbb{N},\sum_{i=1}^{l}r_{i}<r}r_{1}^{v_{1}-1}r_{2}^{v_{2}-1}\cdots r_{l}^{v_{l}-1}\leq C<\infty.$ (9)

### v2

@@
 Fractional Brownian motion; Fourth moment theorem; Berry-Esséen bound; Fractional Ornstein–Uhlenbeck; Sub-bifractional Brownian motion; Fractional Gaussian process; Kolmogorov distance; Malliavin calculus.
-{amscode}
-62M09,60G22,60H10,60H30
+{history}
+[…]
+AMS Subject Classification: 62M09,60G22,60H10,60H30
 1 Introduction and main results
@@
 If the process is observed at discrete time instants $t_{k}=kh,\,k=1,2,\dots,n$ and the estimator $\hat{\theta}_{n}$ is given by (2), then there exists a positive constant $C_{\theta,H,h}$ independent of $n$ such that when $n$ is sufficiently large,
-$\begin{split}d_{Kol}\left(\sqrt{n}(\hat{\theta}_{n}-\theta),\mathcal{N}\right)\leq C_{\theta,H,h}\times\left\{\begin{array}[]{ll}\frac{1}{\sqrt{n}},&\quad\text{if }H\in(0,\,\frac{5}{8}],\\ \frac{1}{n^{3-4H}},&\quad\text{if }H\in(\frac{5}{8},\,\frac{3}{4}),\end{array}\right.\end{split}$ (3)
+[…]
 where the normal random variable $\mathcal{N}\sim N(0,\sigma_{1}^{2})$ with $\sigma_{1}^{2}=\frac{\theta^{2}\sigma_{B}^{2}}{4H^{2}a^{2}}$ and $a=H\Gamma(2H)\theta^{-2H}$, $\sigma_{B}^{2}=2\sum_{j=-\infty}^{+\infty}\rho_{0}^{2}(jh)<+\infty$.
-Remark 1.
+Remark 1.2.
 When $H=\frac{3}{4}$, the scaling before $(\hat{\theta}_{n}-\theta)$ is $\sqrt{\frac{n}{\log n}}$ and the corresponding upper bound is $\frac{1}{\log n}$.
@@
 where the driving Gaussian noise $G_{t}$ satisfies the following hypothesis.
-HYPOTHESIS 1.6.
+HYPOTHESIS 1.7
 For $H\in(0,1)$ and $H\neq\frac{1}{2}$, the covariance function $R(s,t)=\mathbb{E}[G_{t}G_{s}]$ of the centered Gaussian process $(G_{t})_{t\in[0,T]}$ with $G_{0}=0$ satisfies the following three hypotheses:
@@
     holds.
-It is clear that the four Gaussian noises from Example 1.2 to Example 1.5 satisfy Hypothesis 1.6, see [9, 8]. Now we give our second contribution of the present paper.
+It is clear that the four Gaussian noises from Example 1.3 to Example 1.6 satisfy Hypothesis 1.7, see [8, 9]. Now we give our second contribution of the present paper.
 Theorem 1.8.
@@
 $\hat{\theta}_{n}=\left(\frac{1}{{H}\Gamma(2H)n}\sum_{j=1}^{n}Z_{jh}^{2}\right)^{-\frac{1}{2H}}.$ (7)
-If the driving noise satisfies Hypothesis 1.6 with $H\in(0,\frac{1}{2})$, then there exists a positive constant $C_{\theta,H,h}$ independent of $n$ such that when $n$ is sufficiently large,
+If the driving noise satisfies Hypothesis 1.7 with $H\in(0,\frac{1}{2})$, then there exists a positive constant $C_{\theta,H,h}$ independent of $n$ such that when $n$ is sufficiently large,
 $d_{Kol}\left(\sqrt{n}(\hat{\theta}_{n}-\theta),\mathcal{N}\right)\leq C_{\theta,H,h}\times\frac{1}{\sqrt{n}}$ (8)
@@
 The inequality (6) is good enough in some applications, see [8, 9]. However, in some situations, a more steep inequality (9) is needed, see [5]. We write it as a new Hypothesis.
-HYPOTHESIS 1.8.
+HYPOTHESIS 1.9
 For $H\in(0,1)$ and $H\neq\frac{1}{2}$, the covariance function $R(s,t)=\mathbb{E}[G_{t}G_{s}]$ of the centered Gaussian process $(G_{t})_{t\in[0,T]}$ with $G_{0}=0$ satisfies the above ($H_{1}$), ($H_{2}$) and the following:
@@
     holds, where $H^{\prime}\in(\frac{1}{2},1),\,K\in(0,2)$ and $H:=H^{\prime}K\in(0,1)$.
-Clearly, both Example 1.2 and Example 1.5 satisfy Hypothesis 1.8. An additional requirement $H^{\prime}\in(\frac{1}{2},1)$ for both Example 1.4 and Example 1.3 makes Hypothesis 1.8 hold.
+Clearly, both Example 1.3 and Example 1.6 satisfy Hypothesis 1.9. An additional requirement $H^{\prime}\in(\frac{1}{2},1)$ for both Example 1.5 and Example 1.4 makes Hypothesis 1.9 hold.
 Theorem 1.10.
 Assume that both the Ornstein-Uhlenbeck model $\{Z_{t}:t\geq 0\}$ and the estimator $\hat{\theta}_{n}$ are given as in Theorem 1.8.
-If the driving noise satisfies Hypothesis 1.8 with $H\in(\frac{1}{2},\frac{3}{4})$, then there exists a positive constant $C_{\theta,H,h}$ independent of $n$ such that when $n$ is sufficiently large,
-$d_{Kol}\left(\sqrt{n}(\hat{\theta}_{n}-\theta),\mathcal{N}\right)\leq C_{\theta,H,h}\times\left\{\begin{array}[]{ll}\frac{1}{\sqrt{n}},&\quad\text{if }H\in(\frac{1}{2},\,\frac{5}{8}],\\ \frac{1}{n^{3-4H}},&\quad\text{if }H\in(\frac{5}{8},\,\frac{3}{4}),\end{array}\right.$ (10)
+If the driving noise satisfies Hypothesis 1.9 with $H\in(\frac{1}{2},\frac{3}{4})$, then there exists a positive constant $C_{\theta,H,h}$ independent of $n$ such that when $n$ is sufficiently large,
+$d_{Kol}\left(\sqrt{n}(\hat{\theta}_{n}-\theta),\mathcal{N}\right)\leq C_{\theta,H,h}\times\left\{\begin{array}[]{ll}\frac{1}{\sqrt{n}},&\quad\text{if }H\in(\frac{1}{2},\,\frac{5}{8}],\\ \frac{1}{n^{3-4H}},&\quad\text{if }H\in(\frac{5}{8},\,\frac{3}{4}),\end{array}\right$.
+(10)
 where $\mathcal{N}$ is the same as in Theorem 1.1.
-Remark 2.
+Remark 1.11.
 The assumption of $H^{\prime}\in(\frac{1}{2},1)$ in Hypothesis ($H_{3}^{\prime}$) rules out $H^{\prime}\in(0,\frac{1}{2}]$, which is not an essential but only a technical requirement.
@@
 If an additional requirement $H^{\prime}\in(\frac{1}{2},1)$ holds for both the bi-fractional Brownian motion and the sub-bifractional Brownian motion, then there exists a positive constant $C_{\theta,H,h}$ independent of $n$ such that when $n$ is sufficiently large,
-$d_{Kol}\left(\sqrt{n}(\hat{\theta}_{n}-\theta),\mathcal{N}\right)\leq C_{\theta,H,h}\times\left\{\begin{array}[]{ll}\frac{1}{\sqrt{n}},&\quad\text{if }H\in(0,\,\frac{5}{8}],\\ \frac{1}{n^{3-4H}},&\quad\text{if }H\in(\frac{5}{8},\,\frac{3}{4}),\end{array}\right.$ (11)
+$d_{Kol}\left(\sqrt{n}(\hat{\theta}_{n}-\theta),\mathcal{N}\right)\leq C_{\theta,H,h}\times\left\{\begin{array}[]{ll}\frac{1}{\sqrt{n}},&\quad\text{if }H\in(0,\,\frac{5}{8}],\\ \frac{1}{n^{3-4H}},&\quad\text{if }H\in(\frac{5}{8},\,\frac{3}{4}),\end{array}\right$.
+(11)
 where $\mathcal{N}$ is the same as in Theorem 1.1.
@@
 Let $\mathfrak{H}$ be the associated reproducing kernel Hilbert space of the Gaussian process $G$, which is defined as the closure of the space of all real-valued step functions on $[0,T]$, equipped with the inner product
-[…]
+$\langle\mathbbm{1}_{[a,b]},\,\mathbbm{1}_{[c,d]}\rangle_{\mathfrak{H}}=\mathbb{E}\left((G_{b}-G_{a})(G_{d}-G_{c})\right),$
 for any $0\leq a<b\leq T$ and $0\leq c<d\leq T$.
@@
 When $H\in(\frac{1}{2},1)$ or the Lebesgue measure of intersection of the supports about two function $f,g\in\mathfrak{H}$ is zero, Mishura [19] provides that
-[…]
+$\langle g,h\rangle_{\mathfrak{H}_{1}}=H(2H-1)\int_{\mathbb{R}^{2}}g(u)h(v)|u-v|^{2H-2}dudv.$
 Suppose that $\mathcal{V}_{[0,T]}$ is the set of functions of bounded variation in $[0,T]$, and by $\mathcal{B}([0,T])$ the Borel $\sigma$-algebra on $[0,T]$.
@@
 If a random variable $X\geq 0$ almost surely, then there exists a positive constant $C$ independent of $T$ such that
-$d_{Kol}(\sqrt{T}(f(X)-\theta),\,\xi)\leq C\times\left(d_{Kol}(\sqrt{T}(X-\mathbb{E}[X]),\eta)+\sqrt{T}\left|\mathbb{E}[X]-g(\theta)\right|+\frac{1}{\sqrt{T}}\right),$
+$d_{Kol}(\sqrt{T}(f(X)-\theta),\,\xi)\leq C\times\left(d_{Kol}(\sqrt{T}(X-\mathbb{E}[X]),\eta)+\sqrt{T}\left|\mathbb{E}[X]-g(\theta)\right|+\frac{1}{\sqrt{T}}\right),$ (10)
 where the two variance $\sigma^{2}_{2},\,\sigma^{2}_{1}$ satisfy the following relation:
@@
 where the constant $a=g(\theta)=H\Gamma(2H)\theta^{-2H}$.
-Proof.
+Proof 3.2.
 Through standard computations, the fractional Ornstein-Uhlenbeck processes $X_{t}$, known as the solution of (1), admits the explicit representation:
@@
 $\left|\mathbb{E}\left(X_{t}^{2}-Y_{t}^{2}\right)\right|=e^{-\theta t}\left|\mathbb{E}\left(Y_{0}\left(e^{-\theta t}Y_{0}-2Y_{t}\right)\right)\right|\leq 3ae^{-\theta t}.$ (20)
-Substituting this result into (17), we obtain the desired result. ∎
+Substituting this result into (18), we obtain the desired result.
 3.1 The second moment and cumulants of the random variable $W_{n}$
@@
 Let $H\in(0,\frac{3}{4})$ and $W_{n}$ be defined as in (21). When $n$ is large enough, the exist a constant $C$ independent of $n$ such that
-$\begin{split}\left|\mathbb{E}(W_{n}^{2})-\sigma_{B}^{2}\right|\leq C\times\left\{\begin{array}[]{ll}\frac{1}{n},&\quad\text{if }H\in(0,\,\frac{1}{2}],\\ \frac{1}{n^{3-4H}},&\quad\text{if }H\in(\frac{1}{2},\,\frac{3}{4}),\end{array}\right.\end{split}$ (21)
+[…]
 where $\sigma_{B}^{2}$ is a series given by
 $\begin{split}\sigma_{B}^{2}=2\sum_{j=-\infty}^{+\infty}\rho_{0}^{2}(jh)<+\infty.\end{split}$ (23)
-Proof.
+Proof 3.4.
 Firstly, we derive the convergence of above series $\sigma_{B}^{2}$ and denote
@@
 Consequently, we obtain
-$\begin{split}S_{1}&\leq\frac{C}{n}\left[\sum_{j,l=1}^{n}e^{-\theta h(l+j)}+\sum_{j,l=1}^{n}e^{-\theta jh}\left(1+l\right)^{2H-2}\right]\\ &\leq\frac{C}{n}\left[\int_{1}^{\infty}\int_{1}^{\infty}e^{-\theta h(x+y)}\mathrm{d}x\mathrm{d}y+\int_{1}^{\infty}e^{-\theta hx}\mathrm{d}x\int_{1}^{n}y^{2H-2}\mathrm{d}y\right]\\ &\leq\frac{C}{n}\left[1+n^{(2H-1)\vee 0}\right]\leq C\times\left\{\begin{array}[]{ll}\frac{1}{n},&\quad\text{if }H\in(0,\,\frac{1}{2}],\\ \frac{1}{n^{2(1-H)}},&\quad\text{if }H\in(\frac{1}{2},\,1).\end{array}\right.\end{split}$ (33)
+[…]
 For the term $S_{2}$, we firstly have known that if and only if $0<H<\frac{3}{4}$,
 $\begin{split}\sigma_{B}^{2}=2\sum_{j=-\infty}^{+\infty}\rho_{0}^{2}(jh)<\infty.\end{split}$ (35)
-And then, making the change of variable $k=j-l$ yields
+And then, making the change of variable k=jlk=j-l yields
 $\begin{split}\frac{1}{n}\sum_{j,l=1}^{n}\rho_{0}^{2}(\left|j-l\right|h)=\sum_{k=1-n}^{n-1}\rho_{0}^{2}(\left|k\right|h)(1-\frac{|k|}{n})=\sum_{k=1-n}^{n-1}\rho_{0}^{2}(\left|k\right|h)-\frac{1}{n}\sum_{k=1-n}^{n-1}\rho_{0}^{2}(\left|k\right|h)|k|.\end{split}$ (36)
@@
 $\begin{split}\sum_{|k|\geq n}^{\infty}\rho_{0}^{2}(\left|k\right|h)\leq C\sum_{k=n+1}^{\infty}(1+|k|)^{2(2H-2)}\leq C\int_{n}^{\infty}x^{2(2H-2)}\mathrm{d}x\leq Cn^{4H-3},\end{split}$ (38)
-$\begin{split}\frac{4}{n}\sum_{k=1}^{n}k\rho_{0}^{2}(kh)&\leq\frac{C}{n}\sum_{k=1}^{n}(1+k)^{2(2H-2)+1}\leq\frac{C}{n}\int_{n}^{\infty}x^{2(2H-2)+1}\mathrm{d}x\\ &\leq Cn^{(4H-2)\vee 0-1}=C\times\left\{\begin{array}[]{ll}\frac{1}{n},&\quad\text{if }H\in(0,\,\frac{1}{2}],\\ \frac{1}{n^{1-2H}},&\quad\text{if }H\in(\frac{1}{2},\,\frac{3}{4}).\end{array}\right.\end{split}$ (38)
+[…]
 As a result, we get the estimation of $S_{2}$ as
-$\begin{split}S_{2}\leq C\times\left\{\begin{array}[]{ll}\frac{1}{n},&\quad\text{if }H\in(0,\,\frac{1}{2}],\\ \frac{1}{n^{1-2H}},&\quad\text{if }H\in(\frac{1}{2},\,\frac{3}{4}).\end{array}\right.\end{split}$ (39)
-Substituting the estimations (33) and (39) into (28) with the fact that $32(H-1)<4H-$ 3 if $H\in(\frac{1}{2},\,\frac{3}{4})$, we obtain the desired result. ∎
+[…]
+Substituting the estimations (34) and (40) into (29) with the fact that 2(H1)<4H32(H-1)<4H-3 if $H\in(\frac{1}{2},\,\frac{3}{4})$, we obtain the desired result.
 Next, we derive the the upper bounds of the third and fourth cumulants of $W_{n}$.
@@
 Let $H\in(0,\frac{3}{4})$ and $W_{n}$ be defined as in (21). Then for large enough $n$, we have
-$\begin{split}\max\left\{\left|k_{3}(W_{n})\right|,k_{4}(W_{n})\right\}\leq C\times\left\{\begin{array}[]{ll}\frac{1}{\sqrt{n}},&\quad\text{if }H\in(0,\,\frac{2}{3}],\\ n^{\frac{3}{2}(4H-3)},&\quad\text{if }H\in(\frac{2}{3},\,\frac{3}{4}).\end{array}\right.\end{split}$ (40)
-Proof.
+[…]
+Proof 3.6.
 The core approach involves comparing the third and fourth cumulants of $W_{n}$ with those of $\overline{W}_{n}$, which denote by a random variable
@@
 [3] and the inequality (5) in Lemma 5.2, we obtain that
-$\begin{split}k_{3}(\overline{W}_{n})&\leq\frac{C}{\sqrt{n}}\left(\sum_{|k|<n}\left|\rho_{0}(k)\right|^{\frac{3}{2}}\right)^{2}\leq\frac{C}{\sqrt{n}}\left(\sum_{k=0}^{n-1}(1+k)^{\frac{3}{2}(2H-2)}\right)^{2}\\ &\leq Cn^{(6H-4)\vee 0-\frac{1}{2}}=C\times\left\{\begin{array}[]{ll}\frac{1}{\sqrt{n}},&\quad\text{if }H\in(0,\,\frac{2}{3}],\\ n^{\frac{3}{2}(4H-3)},&\quad\text{if }H\in(\frac{2}{3},\,\frac{3}{4}),\end{array}\right.\end{split}$ (44)
-$\begin{split}k_{4}(\overline{W}_{n})&\leq\frac{C}{n}\left(\sum_{|k|<n}\left|\rho_{0}(k)\right|^{\frac{4}{3}}\right)^{3}\leq\frac{C}{n}\left(\sum_{k=0}^{n-1}(1+k)^{\frac{4}{3}(2H-2)}\right)^{3}\\ &\leq Cn^{(8H-5)\vee 0-1}=C\times\left\{\begin{array}[]{ll}\frac{1}{n},&\quad\text{if }H\in(0,\,\frac{5}{8}],\\ n^{2(4H-3)},&\quad\text{if }H\in(\frac{5}{8},\,\frac{3}{4}).\end{array}\right.\end{split}$ (45)
+[…]
+[…]
 On the other hand, from the identity (19), we rewrite $W_{n}$ as
@@
 where $R_{jh}=-2Y_{jh}Y_{0}+e^{-\theta jh}Y_{0}^{2}$, which satisfies
-$\begin{split}\sup_{j}\left|\left|R_{j}\right|\right|_{L^{2}(\Omega)}<\infty,\end{split}$ (47)
+[…]
 based on the stationary property of $Y_{t}$ and Cauchy-Schwarz inequality. Combining this result with the fact that $R_{j}$ is a 2-th Wiener chaos, we have
-$\begin{split}\sup_{j}\left|\left|\frac{1}{\sqrt{n}}\sum_{j=1}^{n}e^{-\theta jh}R_{jh}\right|\right|_{L^{2}(\Omega)}<\frac{C}{\sqrt{n}},\end{split}$ (48)
+[…]
 where C is independent of $n$. Then, by the identity (47) and Cauchy-Schwarz inequality, Minkowski’s inequality, and hypercontractivity property of Wiener chaos, we obtain that
-$\begin{split}&\left|k_{3}({W}_{n})-k_{3}(\overline{W}_{n})\right|=\left|\mathbb{E}\left(W_{n}^{3}-\overline{W}_{n}^{3}\right)\right|\\ &=\left|\mathbb{E}\left[\left(\frac{1}{\sqrt{n}}\sum_{j=1}^{n}e^{-\theta jh}R_{jh}-\sqrt{n}\left(\mathbb{E}(B_{n})-a\right)\right)\left(W_{n}^{2}+W_{n}\overline{W}_{n}+\overline{W}_{n}^{2}\right)\right]\right|\\ &\leq\left[\left|\left|\frac{1}{\sqrt{n}}\sum_{j=1}^{n}e^{-\theta jh}R_{jh}\right|\right|_{L^{2}(\Omega)}+\left|\left|\sqrt{n}\left(\mathbb{E}(B_{n})-a\right)\right|\right|_{L^{2}(\Omega)}\right]\left|\left|W_{n}^{2}+W_{n}\overline{W}_{n}+\overline{W}_{n}^{2}\right|\right|_{L^{2}(\Omega)}\\ &\leq\frac{C}{\sqrt{n}},\end{split}$ (49)
-where in the last inequality we also have used the estimation (48) and Propositions 3.1, 3.2. A similar method yields
+[…]
+where in the last inequality we also have used the estimation (49) and Propositions 3.1, 3.3. A similar method yields
 $\begin{split}&\left|k_{4}({W}_{n})-k_{4}(\overline{W}_{n})\right|=\left|\mathbb{E}\left(W_{n}^{4}\right)-\mathbb{E}\left(\overline{W}_{n}^{4}\right)\right|+3\left|\left(\mathbb{E}W_{n}^{2}\right)^{2}-\left(\mathbb{E}\overline{W}_{n}^{2}\right)^{2}\right|\\ &\leq\left|\left|\frac{1}{\sqrt{n}}\sum_{j=1}^{n}e^{-\theta jh}R_{jh}-\sqrt{n}\left(\mathbb{E}(B_{n})-a\right)\right|\right|_{L^{2}(\Omega)}\cdot\left|\left|W_{n}^{3}+W_{n}^{2}\overline{W}_{n}++W_{n}\overline{W}_{n}^{2}+\overline{W}_{n}^{3}\right|\right|_{L^{2}(\Omega)}\\ &\quad+3\left|\left(\mathbb{E}W_{n}^{2}+\mathbb{E}\overline{W}_{n}^{2}\right)\left(\mathbb{E}W_{n}^{2}-\mathbb{E}\overline{W}_{n}^{2}\right)\right|\\ &\leq\frac{C}{\sqrt{n}}.\end{split}$ (51)
-Combining the estimations (44), (45), (49), (50), we can obtain the desired result. ∎
+Combining the estimations (45), (46), (50), (51), we can obtain the desired result.
 3.2 Berry-Esséen type upper bound for the moment estimator of fOU process
@@
 by the triangle inequality. The optimal fourth moment theorem of Nourdin and Peccati [22] and the well-known fact that $d_{Kol}(\cdot,\cdot)\leq d_{TV}(\cdot,\cdot)$ imply that
-$\begin{split}d_{Kol}(W_{n},\varpi_{n})&\leq d_{TV}(W_{n},\varpi_{n})\leq C\max\left\{k_{3}(W_{n}),k_{4}(W_{n})\right\}\\ &\leq C\times\left\{\begin{array}[]{ll}\frac{1}{\sqrt{n}},&\quad\text{if }H\in(0,\,\frac{2}{3}],\\ n^{\frac{3}{2}(4H-3)},&\quad\text{if }H\in(\frac{2}{3},\,\frac{3}{4}),\end{array}\right.\end{split}$ (54)
+[…]
 where the last inequality is resulted from Proposition 3.5. Using Proposition 3.6.1 of Nourdin and Peccati [21] and Proposition 3.3 yield
-$\begin{split}d_{Kol}(\varpi_{n},\varpi)\leq\frac{2}{\sigma_{n}^{2}\vee\sigma_{B}^{2}}\left|\sigma_{n}^{2}-\sigma_{B}^{2}\right|\leq C\times\left\{\begin{array}[]{ll}\frac{1}{n},&\quad\text{if }H\in(0,\,\frac{1}{2}],\\ \frac{1}{n^{3-4H}},&\quad\text{if }H\in(\frac{1}{2},\,\frac{3}{4}).\end{array}\right.\end{split}$ (55)
+[…]
 Combining this result with the inequalities (54), (55) implies that
-$\begin{split}(d_{Kol}(\sqrt{n}(B_{n}-\mathbb{E}[B_{n}]),\varpi)\leq C\times\left\{\begin{array}[]{ll}\frac{1}{\sqrt{n}},&\quad\text{if }H\in(0,\,\frac{5}{8}],\\ \frac{1}{n^{3-4H}},&\quad\text{if }H\in(\frac{5}{8},\,\frac{3}{4}).\end{array}\right.\end{split}$ (56)
+[…]
 Secondly, it is straightforward to show for $H\in(0,1)$
@@
 Prior to proving Theorem 1.8, liking the definition of $W_{n}$, we first denote by $\widetilde{W}_{n}$ a second Wiener chaos with respect to the general Gaussian process $G_{t}$ as:
-$\widetilde{W}_{n}=\sqrt{n}\left(A_{n}-\mathbb{E}(A_{n})\right)=\frac{1}{\sqrt{n}}\sum_{j=1}^{n}\left(Z_{jh}^{2}-\mathbb{E}\left(Z_{jh}^{2}\right)\right).$ (1)
+$\widetilde{W}_{n}:=\sqrt{n}\left(A_{n}-\mathbb{E}(A_{n})\right)=\frac{1}{\sqrt{n}}\sum_{j=1}^{n}\left(Z_{jh}^{2}-\mathbb{E}\left(Z_{jh}^{2}\right)\right).$ (1)
 Establishing the Berry-Esséen upper bound for $\widetilde{W}_{n}$ constitutes a critical step in the proof of Theorem 1.8.
@@
 where $\sigma_{B}^{2}$ is a series defined in (23).
-Proof.
+Proof 4.2.
 Based on Proposition 3.3, the proof reduces to verifying that
@@
 $\begin{split}\mathbb{E}(\widetilde{W}_{n}^{2}-W_{n}^{2})&=\frac{2}{n}\sum_{j,l=1}^{n}\left[\tilde{\rho}^{2}(jh,lh)-\rho^{2}(jh,lh)\right]\\ &=\frac{2}{n}\sum_{j,l=1}^{n}\left[\left(\tilde{\rho}(jh,lh)-\rho(jh,lh)+\rho(jh,lh)\right)^{2}-\rho^{2}(jh,lh)\right]\\ &=\frac{2}{n}\sum_{j,l=1}^{n}\left(\tilde{\rho}(jh,lh)-\rho(jh,lh)\right)^{2}+\frac{4}{n}\sum_{j,l=1}^{n}\rho(jh,lh)\left(\tilde{\rho}(jh,lh)-\rho(jh,lh)\right)\\ &:=D_{1}+D_{2}.\end{split}$ (4)
-Next, we will estimate the upper bound of $D_{1}$ and $D_{2}$, respectively. The estimation (6) in Lemma 5.2, along with the symmetry of $j,l$ and the change variable $k=l-j$ imply that
+Next, we will estimate the upper bound of $D_{1}$ and $D_{2}$, respectively. The estimation (6) in Lemma 5.2, along with the symmetry of $j,l$ and the change variable k=ljk=l-j imply that
 $\begin{split}0\leq D_{1}&\leq\frac{C}{n}\sum_{j,l=1}^{n}\left[\left(1+(j\wedge l)\right)^{2(H-1)}\wedge\left(1+\left|j-l\right|\right)^{H-1}\right]^{2}\\ &\leq\frac{C}{n}\sum_{1\leq j\leq l\leq n}\left[\left(1+j\right)^{2(H-1)}\wedge\left(1+(l-j)\right)^{H-1}\right]^{2}\\ &\leq\frac{C}{n}\sum_{1\leq j\leq l\leq n}\left(1+j\right)^{2(H-1)}\cdot\left(1+(l-j)\right)^{H-1}\\ &\leq\frac{C}{n}\sum_{j=1}^{n}\left(1+j\right)^{2(H-1)}\sum_{k=0}^{n-1}\left(1+k\right)^{H-1}\\ &\leq\frac{C}{n}n^{(2H-1)\vee 0}n^{H}=Cn^{H-1}\leq C\frac{1}{\sqrt{n}},\end{split}$ (5)
@@
 $\begin{split}\left|D_{2}\right|&\leq\frac{C}{n}\sum_{j,l=1}^{n}\left(1+\left|j-l\right|\right)^{2(H-1)}\cdot\left[\left(1+(j\wedge l)\right)^{2(H-1)}\wedge\left(1+\left|j-l\right|\right)^{H-1}\right]\\ &\leq\frac{C}{n}\sum_{1\leq j\leq l\leq n}\left(1+(l-j)\right)^{2(H-1)}\cdot\left(1+j\right)^{2(H-1)}\\ &\leq\frac{C}{n}\left(\sum_{j=1}^{n}\left(1+j\right)^{2(H-1)}\right)^{2}\leq\frac{C}{n}n^{(4H-2)\vee 0}=\frac{C}{n}.\end{split}$ (6)
-Consequently, substituting inequality (5), (6) into equation (4) obtains the desired result (3). In summary, this completes the proof. ∎
+Consequently, substituting inequality (5), (6) into equation (4) obtains the desired result (3). In summary, this completes the proof.
 Next, we focus on estimating the third and fourth cumulants of random variable $\widetilde{W}_{n}$.
@@
 $\begin{split}\left|k_{3}(\widetilde{W}_{n})\right|\leq C\times\frac{1}{\sqrt{n}}.\end{split}$ (8)
-Proof.
+Proof 4.4.
 According to the estimations (45) and (50) in Proposition 3.5, we only need to show
@@
 $\begin{split}0\leq I\leq Cn.\end{split}$ (19)
-This complete the proof. ∎
+This complete the proof.
 Proposition 4.5.
@@
 $\begin{split}k_{4}(\widetilde{W}_{n})\leq C\times\frac{1}{\sqrt{n}}.\end{split}$ (21)
-Proof.
+Proof 4.6.
 From the previous estimation concerning $k_{4}(\overline{W}_{n})$ and $\left|k_{4}({W}_{n})-k_{4}(\overline{W}_{n})\right|$ in Proposition 3.5, it is sufficient to prove that
@@
 $\begin{split}J\leq\sum_{i=1}^{15}\left|J_{i}\right|\leq Cn^{2H}\leq Cn^{\frac{3}{2}}.\end{split}$ (34)
-This completes the proof. ∎
+This completes the proof.
 4.2 Proofs of main Theorems
@@
 $\begin{split}\sqrt{n}\left|\mathbb{E}(B_{n}-A_{n})\right|&\leq\frac{1}{\sqrt{n}}\sum_{j=1}^{n}\left|\mathbb{E}\left(Z_{jh}^{2}-X_{jh}^{2}\right)\right|\leq\frac{C}{\sqrt{n}}\sum_{j=1}^{n}\left(1\wedge(jh)^{2(H-1)}\right)\\ &\leq\frac{C}{\sqrt{n}}\sum_{j=1}^{n}\left(1+jh\right)^{2(H-1)}\leq C\times n^{(2H-1)\vee 0-\frac{1}{2}}=C\times\frac{1}{\sqrt{n}}.\end{split}$ (37)
-Next, using arguments analogous to those in the estimation of (56) combined with Propositions 4.1, 4.2, 4.3, we derive that for $H\in(0,\frac{1}{2})$,
+Next, using arguments analogous to those in the estimation of (57) combined with Propositions 4.1, 4.3, 4.5, we derive that for $H\in(0,\frac{1}{2})$,
 $\begin{split}d_{Kol}(\sqrt{n}(A_{n}-\mathbb{E}[A_{n}]),\varpi)\leq C\times\frac{1}{\sqrt{n}}.\end{split}$ (38)
@@
 Proof of Theorem 1.10.
-The distinction between Theorem 1.7 and Theorem 1.9 lies in improving Hypothesis 1.6 with $H\in(0,\frac{1}{2})$ to Hypothesis 1.8 with $H\in(\frac{1}{2},\frac{3}{4})$, thereby accommodating broader Gaussian processes such as the bi-fractional Brownian motion and the sub-bifractional Brownian motion.
+The distinction between Theorem 1.8 and Theorem 1.10 lies in improving Hypothesis 1.7 with $H\in(0,\frac{1}{2})$ to Hypothesis 1.9 with $H\in(\frac{1}{2},\frac{3}{4})$, thereby accommodating broader Gaussian processes such as the bi-fractional Brownian motion and the sub-bifractional Brownian motion.
 Notice that the estimations in Lemma 5.2 play a key rule in the proof of Theorem 1.8.
@@
 Proposition 4.7.
-Let ${\rho}(t,s)=\mathbb{E}(X_{t}X_{s})$, $\tilde{\rho}(t,s)=\mathbb{E}(Z_{t}Z_{s})$ be the covariance function of the Ornstein-Uhlenbeck processes $X_{t}$ and $Z_{t}$ driven by fBm $B_{t}^{H}$ and $G_{t}$ satisfying Hypothese 1.8 with $H\in(\frac{1}{2},1)$.
+Let ${\rho}(t,s)=\mathbb{E}(X_{t}X_{s})$, $\tilde{\rho}(t,s)=\mathbb{E}(Z_{t}Z_{s})$ be the covariance function of the Ornstein-Uhlenbeck processes $X_{t}$ and $Z_{t}$ driven by fBm $B_{t}^{H}$ and $G_{t}$ satisfying Hypothese 1.9 with $H\in(\frac{1}{2},1)$.
 Then there exists a constant $C\geq 0$ independent of $T$ such that for any $0\leq s\leq t\leq T$,
@@
 $\left|\mathbb{E}[Z_{t}^{2}]-\mathbb{E}[X_{t}^{2}]\right|\leq C\left(1\wedge t^{2(H-1)}\right).$ (41)
-Proof.
+Proof 4.8.
 For any $0\leq s\leq t\leq T$, according to the relationship (6) between the inner products of two functions in the Hilbert spaces $\mathfrak{H}$ and $\mathfrak{H}_{1}$, we have
 $\left|\mathbb{E}[Z_{t}^{2}]-\mathbb{E}[X_{t}^{2}]\right|\leq\int_{0}^{t}e^{-\theta(t-u)}\mathrm{d}u\int_{0}^{t}e^{-\theta(t-v)}\left|\frac{\partial^{2}R(u,v)}{\partial u\partial v}-\frac{\partial^{2}R^{B}(u,v)}{\partial u\partial v}\right|\mathrm{d}v.$ (42)
-At the same time, Hypothesis 1.8 with restriction $H\in(\frac{1}{2},1)$ implies that
+At the same time, Hypothesis 1.9 with restriction $H\in(\frac{1}{2},1)$ implies that
 $\left|\frac{\partial}{\partial s}\left(\frac{\partial R(s,t)}{\partial t}-\frac{\partial R^{B}(s,t)}{\partial t}\right)\right|\leq C_{1}(t+s)^{2H-2}+C_{2}(s^{2H^{\prime}}+t^{2H^{\prime}})^{K-2}(st)^{2H^{\prime}-1},$ (43)
@@
 $\begin{split}{\rm II}&:=\int_{0}^{s}e^{-\theta(s-u)}\mathrm{d}u\int_{s}^{t}e^{-\theta(t-v)}\left|\frac{\partial^{2}R(u,v)}{\partial u\partial v}-\frac{\partial^{2}R^{B}(u,v)}{\partial u\partial v}\right|\mathrm{d}v\\ &\leq C_{1}\int_{0}^{s}e^{-\theta(s-u)}\mathrm{d}u\int_{s}^{t}e^{-\theta(t-v)}(u+v)^{2H-2}\mathrm{d}v\\ &+C_{2}\int_{0}^{s}e^{-\theta(s-u)}\mathrm{d}u\int_{s}^{t}e^{-\theta(t-v)}(u^{2H^{\prime}}+v^{2H^{\prime}})^{K-2}(uv)^{2H^{\prime}-1}\mathrm{d}v\\ &:={\rm II}_{1}+{\rm II}_{2},\end{split}$ (46)
-under Hypothesis 1.8. Let’s estimate ${\rm II}$ in two parts based on the inequality (43).
-Part 1. Making the change of variable $x=v-s$ implies that
+under Hypothesis 1.9. Let’s estimate $\rm II$} in two parts based on the inequality (43).
+Part 1. Making the change of variable x=vsx=v-s implies that
 $\begin{split}{\rm II}_{1}\leq C\int_{0}^{s}e^{-\theta(s-u)}\mathrm{d}u\int_{0}^{t-s}e^{-\theta((t-s)-x)}x^{2H-2}\mathrm{d}x\leq C(t-s)^{2H-2}\end{split}$ (47)
-where in the last inequality we use Lemma 6 with $H\in(\frac{1}{2},1)$.
-Part 2. Suppose that $H^{\prime}\in(\frac{1}{2},1),\,K\in(0,2)$ and $H:=H^{\prime}K\in(\frac{1}{2},1)$. We make the change of variable $x=v-s$ and use the L’Hôpital’s rule to get that
+where in the last inequality we use Lemma 5.1 with $H\in(\frac{1}{2},1)$.
+Part 2. Suppose that $H^{\prime}\in(\frac{1}{2},1),\,K\in(0,2)$ and $H:=H^{\prime}K\in(\frac{1}{2},1)$. We make the change of variable x=vsx=v-s and use the L’Hôpital’s rule to get that
 $\begin{split}&\lim_{y\rightarrow\infty}\frac{{\rm II}_{2}}{y^{2(H^{\prime}K-1)}}\\ &\leq C\lim_{y\rightarrow\infty}\frac{\int_{0}^{s}e^{-\theta(s-u)}u^{2H^{\prime}-1}\mathrm{d}u\int_{0}^{y}e^{\theta x}(u^{2H^{\prime}}+(x+s)^{2H^{\prime}})^{K-2}(x+s)^{2H^{\prime}-1}\mathrm{d}x}{y^{2(H^{\prime}K-1)}e^{\theta y}}\\ &=C\lim_{y\rightarrow\infty}\frac{\int_{0}^{s}e^{-\theta(s-u)}u^{2H^{\prime}-1}\mathrm{d}u(u^{2H^{\prime}}+(y+s)^{2H^{\prime}})^{K-2}(y+s)^{2H^{\prime}-1}}{\theta y^{2(H^{\prime}K-1)}+2(H^{\prime}K-1)y^{2(H^{\prime}K-1)-1}}\\ &=C\lim_{y\rightarrow\infty}\int_{0}^{s}e^{-\theta(s-u)}(\frac{u}{y})^{2H^{\prime}-1}\left[(\frac{u}{y})^{2H^{\prime}}+(1+\frac{s}{y})^{2H^{\prime}}\right]^{K-2}\mathrm{d}u\frac{(1+\frac{s}{y})^{2H^{\prime}-1}}{\theta+2(H^{\prime}K-1)y^{-1}},\end{split}$ (48)
@@
 ${\rho}(t,s)=\mathbb{E}(X_{t}X_{s}),\quad\tilde{\rho}(t,s)=\mathbb{E}(Z_{t}Z_{s})$
-by the covariance function of the Ornstein-Uhlenbeck processes $X_{t}$ and $Z_{t}$ driven by fBm $B_{t}^{H}$ and $G_{t}$ satisfying Hypothese 1.6.
+by the covariance function of the Ornstein-Uhlenbeck processes $X_{t}$ and $Z_{t}$ driven by fBm $B_{t}^{H}$ and $G_{t}$ satisfying Hypothese 1.7.
 Then there exists a positive constant $C$ independent of $T$ such that for any $0\leq s\leq t\leq T$,
@@
 $\sum_{r_{i}\in\mathbb{N},\sum_{i=1}^{l}r_{i}<r}r_{1}^{v_{1}-1}r_{2}^{v_{2}-1}\cdots r_{l}^{v_{l}-1}\leq C<\infty.$ (9)
+Next, we provide a detailed proof of the key Lemma 2.1 of this paper.
+Proof of Lemma 2.1.
+Without loss of the generality, we assume $\sigma_{1}=1$ and generalize the function $f(x)=x^{-\frac{1}{\alpha}},\,0<\alpha<2$.
+Then we have $\xi\sim N(0,1)$ and $\Phi(z)$ is its cumulative distribution function.
+Thus, we have
+$\begin{split}d_{Kol}(\sqrt{T}(f(X)-\theta),\,\xi)=\sup_{z\in\mathbb{R}}\left|P\left({\sqrt{T}}(f(X)-\theta)\leq z\right)-\Phi(z)\right|\end{split}$ (10)
+Denote
+$\begin{split}A(z):=P\left({\sqrt{T}}(f(X)-\theta)\leq z\right)-\Phi(z).\end{split}$ (11)
+Since $X\geq 0$ almost surely and $f(x)=x^{-\frac{1}{\alpha}},\,0<\alpha<2$, we have $f(X)>0$ a.s.
+Hence, we shall assume $z>-\sqrt{T}\theta$.
+Otherwise, the standard estimate for a normal random variable $\Phi(-t)\leq\frac{1}{2t},\,\forall t>0$ yields
+$|A|=\Phi(z)\leq\Phi\big(-\sqrt{T}\theta\big)\leq\frac{C}{\sqrt{T}}\,.$
+When $z>-\sqrt{T}\theta$, we have $\frac{z}{\sqrt{T}}+\theta>0$. Hence, the monotonicity of the function $f(x)=x^{-\frac{1}{\alpha}}$ implies that
+$\left\{X^{-\frac{1}{\alpha}}-\theta\leq\frac{z}{\sqrt{T}}\right\}=\left\{X-\theta^{-\alpha}\geq\left(\frac{z}{\sqrt{T}}+\theta\right)^{-\alpha}-\theta^{-\alpha}\right\}.$
+Recall that $g(\theta)=\theta^{-\alpha}$. We have when $z>-\sqrt{T}\theta$,
+$\begin{split}A(z)&=P\left(X^{-\frac{1}{\alpha}}-\theta\leq\frac{z}{\sqrt{T}}\right)-\Phi(z)\\ &=P\left(X-\theta^{-\alpha}\geq\left(\frac{z}{\sqrt{T}}+\theta\right)^{-\alpha}-\theta^{-\alpha}\right)-\Phi(z)\\ &=P\left(\sqrt{\frac{T}{\sigma_{2}^{2}}}\left(X-g(\theta)\right)\geq\sqrt{\frac{T}{\sigma_{2}^{2}}}\theta^{-\alpha}\left[\left(\frac{z}{\sqrt{T}\theta}+1\right)^{-\alpha}-1\right]\right)-\Phi(z)\\ &=P\left(\sqrt{\frac{T}{\sigma_{2}^{2}}}\left(X-g(\theta)\right)\geq\frac{\sqrt{T}\theta}{\alpha}\left[\left(\frac{z}{\sqrt{T}\theta}+1\right)^{-\alpha}-1\right]\right)-\Phi(z),\end{split}$
+where in the last line, we use (11) and $g^{\prime}(\theta)=-\alpha\theta^{-\alpha-1}$. Next, we take the short-hand notion $\bar{\Phi}(z)=1-\Phi(z)$ and
+$\nu=\frac{\sqrt{T}\theta}{\alpha}\left[\left(\frac{z}{\sqrt{T}\theta}+1\right)^{-\alpha}-1\right].$
+It is clear that
+$\left\{\sqrt{\frac{T}{\sigma_{2}^{2}}}\left(X-g(\theta)\right)\geq\nu\right\}=\left\{\sqrt{\frac{T}{\sigma_{2}^{2}}}\left(X-\mathbb{E}[X]\right)\geq\nu-\sqrt{\frac{T}{\sigma_{2}^{2}}}\left(\mathbb{E}[X]-g(\theta)\right)\right\}.$
+Denote $\bar{\Phi}(z)=1-\Phi(z)$. Hence, the triangle inequality implies that
+$\begin{split}\left|A(z)\right|&\leq\left|P\left(\sqrt{\frac{T}{\sigma_{2}^{2}}}\left(X-\mathbb{E}[X]\right)\geq\nu-\sqrt{\frac{T}{\sigma_{2}^{2}}}\left(\mathbb{E}[X]-g(\theta)\right)\right)-\bar{\Phi}\left(\nu-\sqrt{\frac{T}{\sigma_{2}^{2}}}\left(\mathbb{E}[X]-g(\theta)\right)\right)\right|\\ &\quad+\left|\bar{\Phi}\left(\nu-\sqrt{\frac{T}{\sigma_{2}^{2}}}\left(\mathbb{E}[X]-g(\theta)\right)\right)-\bar{\Phi}\left(\nu\right)\right|+\left|\bar{\Phi}\left(\nu\right)-\Phi(z)\right|\\ &\leq d_{Kol}(\sqrt{\frac{T}{\sigma_{2}^{2}}}\left(X-\mathbb{E}[X]\right),\,\xi)+\sqrt{\frac{T}{\sigma_{2}^{2}}}\left|\mathbb{E}[X]-g(\theta)\right|+\left|\bar{\Phi}\left(\nu\right)-\Phi(z)\right|\\ &\leq C\Bigg(d_{Kol}(\sqrt{T}(X-\mathbb{E}[X]),\eta)+\sqrt{T}\left|\mathbb{E}[X]-g(\theta)\right|+\frac{1}{\sqrt{T}}\Bigg),\end{split}$ (12)
+where in the last two inequalities we use the standard estimate for the tail of a normal random variable, $\left|\bar{\Phi}(z_{1})-\bar{\Phi}(z_{2})\right|\leq\left|z_{1}-z_{2}\right|$, and Lemma 5.4 of [9].
+Subscribing the inequality (12) into (10) yields (10).
+The following it taken from Lemma 5.4 of [9].
+Lemma 5.4.
+Let $c>0$ be a constant.
+Denote $\nu(z)=\frac{c}{2\beta}\sqrt{T}\Big[\big(1+\frac{z}{c\sqrt{T}}\big)^{-2\beta}-1\Big],\,0<\beta<1$, when $z>-c\sqrt{T}$ and $\bar{\Phi}(z)=1-P(Z\leq z)$.
+Then there exists some positive number $C$ independent of $T$ such that
+$\sup_{z>-c\sqrt{T}}\left|\bar{\Phi}(\nu)-\Phi(z)\right|\leq\frac{C}{\sqrt{T}}.$
+Proof 5.5.
+We follow the line of the proof of Theorem 3.2 in [24]. By the mean value theorem, there exists some number $\eta(z)\in(0,1)$ such that
+$\nu=-z\big(1+\frac{z\eta}{c\sqrt{T}}\big)^{-2\beta-1}.$
+Hence,
+$\left|\bar{\Phi}(\nu)-\Phi(z)\right|$ $=\left|{\Phi}\Big(\big(1+\frac{z\eta}{c\sqrt{T}}\big)^{-2\beta-1}\cdot z\Big)-\Phi(z)\right|$
+$=\frac{1}{\sqrt{2\pi}}{\int^{z}_{z\big(1+\frac{z\eta}{c\sqrt{T}}\big)^{-2\beta-1}}\,e^{-\frac{t^{2}}{2}}\mathrm{d}t}\,.$
+When $z\in(-c\sqrt{T},-\frac{1}{2}c\sqrt{T}]$, it is obvious that
+$\frac{1}{\sqrt{2\pi}}{\int_{z\big(1+\frac{z\eta}{c\sqrt{T}}\big)^{-2\beta-1}}^{z}\,e^{-\frac{t^{2}}{2}}\mathrm{d}t}\leq\Phi(-\frac{1}{2}c\sqrt{T})\leq\frac{C}{\sqrt{T}}.$
+When $z\in(-\frac{1}{2}c\sqrt{T},\,0]$, we have
+$\frac{1}{\sqrt{2\pi}}{\int_{z\big(1+\frac{z\eta}{c\sqrt{T}}\big)^{-2\beta-1}}^{z}\,e^{-\frac{t^{2}}{2}}\mathrm{d}t}$ $\leq\left|z\right|e^{-\frac{z^{2}}{2}}\Big(\big(1+\frac{z\eta}{c\sqrt{T}}\big)^{-2\beta-1}-1\Big).$ (13)
+The mean value theorem implies that there exists some number $\eta^{\prime}\in(0,1)$ such that
+$\big(1+\frac{z\eta}{c\sqrt{T}}\big)^{-2\beta-1}-1=(-1-2\beta)\frac{z\eta}{c\sqrt{T}}\big(1+\frac{z\eta\eta^{\prime}}{c\sqrt{T}}\big)^{-2\beta-2}\leq(1+2\beta)\left(\frac{1}{2}\right)^{-2\beta-2}\frac{\left|z\right|}{c\sqrt{T}}.$
+Substituting the above inequality into (13), and since the function $f(z)={z}^{2}e^{-\frac{z^{2}}{2}}$ is uniformly bounded, we have that when $z\in(-\frac{1}{2}c\sqrt{T},\,0]$,
+$\frac{1}{\sqrt{2\pi}}{\int_{z\big(1+\frac{z\eta}{c\sqrt{T}}\big)^{-2\beta-1}}^{z}\,e^{-\frac{t^{2}}{2}}\mathrm{d}t}$ $\leq\frac{C}{\sqrt{T}}.$
+Denote $a=\sup\left\{z^{2}e^{-\frac{z^{2}}{2}}:\,z\in\mathbb{R}\right\}$ and denote the domain
+$D:=\left\{(s,z)\in\mathbb{R}_{+}^{2}:\,\frac{1}{z}\big(1+\frac{z\eta}{c\sqrt{T}}\big)^{-2\beta-1}\leq s\leq\frac{1}{z},\,z\in(0,\,c\sqrt{T})\right\}.$
+We claim that the function
+$f_{2}(s,z)=z^{2}e^{-\frac{s^{2}z^{4}}{2}},\quad(s,z)\in D$
+is uniformly bounded in the domain $D$. In fact, for the above positive constant $a$, we have
+$f_{2}(s,z)=\frac{1}{s^{2}z^{2}}\left((sz^{2})^{2}e^{-\frac{(sz^{2})^{2}}{2}}\right)\leq\frac{a}{(sz)^{2}}\leq a\big(1+\frac{z\eta}{c\sqrt{T}}\big)^{2(2\beta+1)}\leq 2^{2(2\beta+1)}a.$
+When $z\in(0,\,c\sqrt{T})$, using the mean value theorem and making the change of variable $t=z^{2}s$ together with the fact that $f_{2}(s,z)$ is uniformly bounded in the above domain $D$, we conclude that there exists a number $\eta^{\prime}\in(0,1)$ such that
+${\int_{z\big(1+\frac{z\eta}{c\sqrt{T}}\big)^{-2\beta-1}}^{z}\,e^{-\frac{t^{2}}{2}}\mathrm{d}t}$ $={\int_{\frac{1}{z}\big(1+\frac{z\eta}{c\sqrt{T}}\big)^{-2\beta-1}}^{\frac{1}{z}}\,\,z^{2}e^{-\frac{s^{2}z^{4}}{2}}\mathrm{d}s}$
+$\leq C\frac{1}{z}\Big(1-\big(1+\frac{z\eta}{c\sqrt{T}}\big)^{-2\beta-1}\Big)$
+$=C(1+2\beta)\frac{1}{z}\big(1+\frac{z\eta\eta^{\prime}}{c\sqrt{T}}\big)^{-2\beta-2}\frac{z\eta}{c\sqrt{T}}$
+$\leq\frac{C}{\sqrt{T}}.$
+Finally, when $z\in[c\sqrt{T},\,\infty)$, we have that
+${\Phi}(-z)\leq{\Phi}(-c\sqrt{T})\leq\frac{1}{2c\sqrt{T}}$
+and $\nu(z)\leq\frac{c}{2\beta}\sqrt{T}\Big[2^{-2\beta}-1\Big]$. Hence,
+$\Phi(\nu)$ $\leq\Phi(\frac{c}{2\beta}\sqrt{T}[2^{-2\beta}-1])\leq\frac{\beta}{[1-2^{-2\beta}]c\sqrt{T}},$
+$\left|\bar{\Phi}(\nu)-\Phi(z)\right|$ $=\left|\bar{\Phi}(z)-\Phi(v)\right|\leq\bar{\Phi}(z)+\Phi(\nu)={\Phi}(-z)+\Phi(\nu)\leq\frac{C}{\sqrt{T}}.$
+Acknowledgments
+The authors wish to express their sincere thanks to the anonymous referees for their valuable suggestions and comments.
+Y.
+Chen is supported by the National Natural Science Foundation of China (No.
+12461029); Z.
+Tang are supported by the Project of Basic Research of Science and Technology Plan of Gansu Province in 2025–Excellent doctoral program (25JRRA235) and the “Innovation Star” project of university graduate students in Gansu Province (2025CXZX-867); Y.
+Li are supported by the Youth Project of Hunan Provincial Natural Science Foundation of China (No.
+2024JJ6417), the Project of Education Department of Hunan Province (24B0187) and the 111 Project (No.
+D23017); H.
+Yi is supported by the High level Talent Research Project of Liupanshui Normal University (LPSSYKYJJ202416).